### v1

# Entangled in Representations: Mechanistic Investigation of Cultural Biases in Large Language Models

Abstract
The growing deployment of large language models (LLMs) across diverse cultural contexts necessitates a better understanding of how the overgeneralization of less documented cultures within LLMs’ representations impacts their cultural understanding. Prior work only performs extrinsic evaluation of LLMs’ cultural competence, without accounting for how LLMs’ internal mechanisms lead to cultural (mis)representation. To bridge this gap, we propose Culturescope, the first mechanistic interpretability-based method that probes the internal representations of LLMs to elicit the underlying cultural knowledge space. CultureScope utilizes a patching method to extract the cultural knowledge. We introduce a cultural flattening score as a measure of the intrinsic cultural biases. Additionally, we study how LLMs internalize Western-dominance bias and cultural flattening, which allows us to trace how cultural biases emerge within LLMs. Our experimental results reveal that LLMs encode Western-dominance bias and cultural flattening in their cultural knowledge space. We find that low-resource cultures are less susceptible to cultural biases, likely due to their limited training resources. Our work provides a foundation for future research on mitigating cultural biases and enhancing LLMs’ cultural understanding. Our codes and data used for experiments are publicly available [^1] https://github.com/copenlu/CultureScope.
1 Introduction
Large language models (LLMs) are increasingly being used in culturally diverse contexts, where understanding and responding appropriately to each culture is essential (Pandya and Holia, 2023; Salemi et al., 2023; Liu et al., 2025). However, the cultural knowledge acquired by LLMs is largely shaped by the data they are trained on, which is predominantly Western-centric (Santurkar et al., 2023). This results in severe geographic imbalances, where some regions receive disproportionate attention while others remain marginalized (Nguyen et al., 2023a). While LLMs do acquire some cross-cultural patterns during training (Hershcovich et al., 2022; Arora et al., 2023), this learning often results in overgeneralization. For example, when asked about a popular leisure activity for retired men in Azerbaijan, GPT-4 responded with chess—a plausible answer, but one that reflects general stereotypes of post-Soviet or Eastern European regions rather than Azerbaijan specifically (Myung et al., 2024).
Figure 1: Given the question about the popular domestic vacation spot in Greece, Culturescope first generates an answer to cultural Question at the Inference stage. Then, it reads the hidden representation from the Inference stage and elicits the cultural knowledge used for ‘Zakynthos’ at the Scoping-in stage. We finalize a list of cultural knowledge after the Filtering stage. Culturescope unveils the internal mechanism of LLMs that cannot be revealed through the Inference stage alone.
However, these patterns of overgeneralization directly perpetuate cultural biases such as cultural flattening. Measuring such overgeneralizations through model outputs alone proves insufficient since generated outputs mask the underlying mechanisms that lead to cultural (mis)representation. To facilitate the examination of the underlying mechanisms, we propose to study LLMs’ cultural understanding with mechanistic interpretability (MI) techniques. MI techniques provide us with methods that can directly examine how cultural biases discovered by the extrinsic evaluation Santurkar et al. (2023) is internally processed within model representations, revealing where and how harmful generalizations emerge. We are the first to propose an approach for intrinsic cultural bias evaluation.
In this work, we introduce Culturescope, a method to probe internal representations and surface the cultural knowledge activated during cultural understanding tasks (§4.1). By surfacing this knowledge, we reveal the internal cultural knowledge space from which model outputs are generated. Figure 1 illustrates an overview of Culturescope. To examine the intrinsic ‘cultural flattening’ embedded in the parameter space, we introduce a cultural flattening (CF) score, which quantifies the degree of intersection between cultural knowledge decoded by Culturescope (§4.2). We implement our framework on two cultural understanding tasks, cultural commonsense Question Answering (QA) and extractive QA, across three different LLMs.
We further challenge the model’s cultural understanding by creating multiple-choice questions (MCQs) with hard negatives (§3.2.3). Culturally nuanced answers from high-resource cultures or geographically proximate countries are selected as hard negative options to simulate the Western-dominance bias and the cultural flattening. This setup refrains LLMs to leverage surface-level elimination strategies based on the overgeneralization Khan et al. (2025). Analyzing selected options by LLMs with the attention map method Yuksekgonul et al. (2024) allows us to examine whether extrinsic and intrinsic cultural biases align, by revealing which options the model internally attends to (§5.1).
High CF scores of Western and high-resource cultures show that LLMs encode Western-dominance bias and cultural flattening in their cultural knowledge space (§6.2). Model performances on MCQs with hard negatives demonstrate that low-resource cultures are less likely to be affected by the biases, likely due to their limited training resource (§6.3). This implies that LLMs struggle with low-resource cultures due to the lack of parametric knowledge. In §6.4, we find that LLMs over-attend to tokens from Western and high-resource cultures. This suggests Western-dominance bias is more internalized than the cultural flattening. Our work provides useful signal to future research on mitigating the internalized patterns for biases to build a better culturally aligned model.
2 Related Work
Evaluating Cultural Understanding of LLMs
Previous work has proposed evaluation datasets and frameworks to assess LLMs’ cultural understanding ability acquired during pre-training Keleg and Magdy (2023); Naous and Xu (2025); Pawar et al. (2025). BLEnD Myung et al. (2024) provides a multilingual commonsense QA dataset spanning 16 countries and regions, designed to uncover cross-cultural disparities in everyday knowledge. CAMeL Naous et al. (2024) compares LLM behavior in Arabic versus Western settings across tasks like story generation, NER, and sentiment analysis, exposing systematic cultural biases in LLMs. Other multilingual benchmarks Zhou et al. (2025); Hasan et al. (2025); Wang et al. (2024); Cao et al. (2024) construct culturally localized evaluation datasets that span domains such as cuisine, proverbs, news, and reasoning. Across these datasets, performance gaps are consistently observed between high-resource and underrepresented languages and cultures, often linked to pre-training data imbalances that favor dominant regions Naous and Xu (2025).
While these efforts highlight important cross-cultural disparities, they perform an extrinsic evaluation, overlooking the underlying mechanism and cultural knowledge space embedded in LLMs. To address this gap, our paper aims to reveal how culture is embedded, entangled, or flattened within the models’ inner representations.
Mechanistic Interpretability
MI techniques are developed to explain the inner workings of LLMs by identifying responsible model components, such as neurons and attention heads Meng et al. (2022); Geva et al. (2023); Yu et al. (2024). Leveraging their transparency, recent studies have employed MI techniques to investigate how specific behaviors emerge in LLMs. For instance, they have been used to uncover and manipulate components associated with social biases, enabling both diagnostic and steering interventions Liu et al. (2024); Durmus et al. (2024); Yang et al. (2024). Despite growing interest in the cultural capabilities of LLMs, no prior work has explored cultural understanding through the lens of MI. Our study fills this gap by applying MI techniques to probe the internal representation of cultural knowledge in LLMs, offering new insights into how cultural understanding is encoded and organized within the model.
3 Experimental Setup
3.1 Preliminaries
A dataset $D=[(q_{1},C_{1},a_{1}),...,(q_{N},C_{N},a_{N})]$ consists of $N$ tuple instances containing: a question $q$, an option list $C=[c_{0},c_{1},c_{2},a]$ containing three options for MCQ, one gold answer $a$ and a country of interest $y$. For MCQ, an LLM is given $C$ and $q$ to generate an output $O=[o_{0},...,o_{P}]$ consisting of $P$ tokens. For open-ended QA, an LLM is only given $q$ to generate an output $O$.
To generate an answer, an LLM converts a tokenized input text $T=[t_{0},...,t_{S}]$ containing $S$ number of tokens into $d$-dimenstional vectors using the embedding matrix $E\in\mathbb{R}^{|\mathbb{V}\times d|}$. Then, the vectors are processed through $L$ layers, each containing a multi-head self-attention (MHSA) layer and an MLP layer. The hidden representation $x_{i}^{l}$ from a layer $l$, on a token $t_{i}$ is computed by:
$x_{i}^{l}=x_{i}^{l-1}+a_{i}^{l}+m_{i}^{l}$ (1)
where $a_{i}^{l}$ is an output from the MHSA layer and $m_{i}^{l}$ from the MLP layer. The hidden representation from the last layer $x_{i}^{L}$ is converted into a token by calculating the logits with the unembedding layer.
3.2 Datasets
3.2.1 Cultural QA Datasets
We select BLEnD Myung et al. (2024), a cultural commonsense QA dataset, and CAMeL-2 Naous and Xu (2025), an extractive QA dataset featuring culturally grounded entities. BLEnD Myung et al. (2024) is a hand-crafted benchmark designed to evaluate LLMs’ everyday knowledge across diverse cultures in English. It comprises 500 short-answer question-answer pairs for each country, where the answers vary depending on the country’s cultural or regional context. To reduce computational cost, we exclude North Korea and West Java, resulting in a final selection of 14 countries from the BLEnD dataset.
CAMeL-2 Naous and Xu (2025) is a bilingual benchmark originally constructed to evaluate LLMs’ entity extraction capabilities on Arabic and English entities. With CAMeL-2, an LLM is asked to extract an entity from a context collected from Arabic X/Twitter data according to the specified entity type in the input. We take the English partition and reduce the dataset to 14 countries to keep a similar country distribution to the BLEnD dataset. Dataset details, including domains and countries covered by the datasets, can be found in Appendix A.
3.2.2 Grouping of Cultures
We categorize 14 countries from each dataset along the resource dimension and the region dimension to study how overgeneralization manifests across these dimensions. For the resource dimension, we adopt the taxonomy proposed by Joshi et al. (2020), which classifies languages into six levels (0: very low-resource to 5: very high-resource). For our experiments, we simplify this into three groups: High (Level 5), Mid (Levels 3-4), and Low (Levels 0-2). We assign each country a language resource level based on its most widely spoken language provided by Wikipedia. For the region dimension, we group countries into six regions based on continents. We split Asia into three subregions, which leaves us six regional groups: North America, Europe, Africa, West Asia, South Asia, and East Asia. A complete list of countries within each group is provided in Appendix A.1.
3.2.3 Cultural MCQ with hard negatives
Khan et al. (2025) found that if MCQs lack the adversarial depth to probe genuine cultural understanding, models can exploit surface-level elimination strategies without truly understanding cultural distinctions. Thus, we propose a cultural MCQ with hard negatives to study how overgeneralization—driven by regional or resource dominance or similarity—affects the downstream task. Since BLEnD Myung et al. (2024) provides different answers from each culture with the same question, we create BLEnD-resource and BLEnD-region partition using culturally nuanced answers in BLEnD.
Figure 2: A cultural question about the popular family game in China from the BLEnD dataset Myung et al. (2024). For the given question about China, if an LLM answers ‘Yutnori’, a popular family game in South Korea, it is caused by the effect of cultural flattening between South Korea and China. On the other hand, if the answer is ‘Monopoly’, the LLM is generating an answer from a high resource culture.
We design two types of multiple choice question option lists that incorporate hard negative options: $C_{resource}$ and $C_{region}$, corresponding to BLEnD-resource and BLEnD-region, respectively. For $C_{resource}$, given a question $q$ targeting country $y$, we sample one country from each of the three resource levels excluding $y$. We obtain these three country’s respective gold answers when substituted into $q$ for $y$, resulting in three hard negative options: $c_{high}$, $c_{mid}$, and $c_{low}$. For $C_{region}$, we sample one country from the same geographical region as $y$ (excluding $y$) and extract its corresponding answer to construct a region-based hard negative option, $c_{flatten}$. Two additional options are randomly selected from countries in different regions. All options are shuffled to avoid positional bias. Figure 2 shows the example of $C_{resource}$ and $C_{region}$ for the question about China’s popular family game. Cultural MCQ with hard negative options allows us to examine when a model generates incorrect answers to cultural questions, whether a model’s cultural confusion arises from similarity in resource level or regional proximity.
3.3 Models
Application of MI methods requires a full-access to the model weights. Due to the requirement, we conduct our experiments with three recent open-sourced LLMs: Meta-Llama-3.1-8B-Instruct (Llama-3.1, Grattafiori et al. (2024)), aya-expanse-8b (aya-expanse, Dang et al. (2024)), and Qwen2.5-7B-Instruct (Qwen2.5, Qwen Team (2024)).
3.4 Patchscope
Patchscope Ghandeharioun et al. (2024) utilizes an LLM itself to generate natural language explanations of its internal representations. It consists of two forward passes, one using a source prompt and the other using an inspection prompt, with a patching operation between them. An inspection prompt is designed to guide an LLM as a probe to extract specific knowledge encoded in its internal representations, aligned with a predefined objective, such as next-token prediction, or attribute extraction. Utilizing an LLM itself with an inspection prompt as a probing mechanism addresses key limitations of prior methods Hernandez et al. (2024); Geva et al. (2022); Belrose et al. (2025), which often rely on predefined probing classes or suffer from limited interpretability due to sub-word tokenization.
In the context of cultural knowledge, these limitations are particularly pronounced. It is challenging to exhaustively define all relevant cultural knowledge. Additionally, a direct projection onto the vocabulary space via the unembedding matrix becomes difficult to interpret since cultural knowledge is frequently tokenized into multiple tokens Naous and Xu (2025). To address these challenges, we introduce a Patchscope-based method tailored for probing the cultural knowledge space. To our knowledge, this is the first work to apply interpretability techniques for investigating cultural knowledge in LLMs.
4 Probing Cultural Knowledge within Internal Layers
Probing the cultural knowledge processed by each layer for the given input provides insights into how cultural knowledge for one culture is overlapping with different cultures within the inner layers of an LLM. To translate internal representations of LLMs to natural language that reveals the cultural knowledge space, we propose Culturescope, building upon the existing interpretability method, Patchscope Ghandeharioun et al. (2024).
4.1 Culturescope
Culturescope consists of three stages: inference, scoping-in, and filtering. Our Culturescope allows us to move beyond what is observable from model responses alone, overcoming the limitation of extrinsic evaluation.
Step 1. Inference
An LLM first encodes a tokenized input $T_{i}$ of $i$-th instance and generates an output $O_{i}$, which is an LLM answer to an Open-Ended cultural QA. Then, we compute the representative hidden representations used to generate an answer, which are patched onto the inspection prompt during the scoping-in stage.
As Patchscope does not consider patching with multi-tokens Ghandeharioun et al. (2024), we adopt Bronzini et al. (2024)’s approach originally developed for fact-checking claims to condense an LLM’s cultural answer involving multiple tokens into a single hidden representation. Similarly, we compute the representative hidden representation $x_{*}^{l}$ of the $O_{i}$, which consists of multiple tokens, on the layer $l$. Specifically, we perform the weighted sum of hidden states as in Eq. 2 for the layer $l$. We set the weight $w_{p}$ of each token to one if it is a noun or a verb. Other token weights are set to zero. We use this $x_{*}^{l}$ in the next stage.
$x_{*}^{l}=\sum_{p=0}^{P}x_{p}^{l}*w_{p}$ (2)
Step 2. Scoping-in
At this stage, we further elicit the cultural knowledge encoded in $x_{*}^{l}$ to reveal the cultural knowledge space utilized for the $O_{i}$. To elicit a list of cultural knowledge $\mathcal{CK}_{i}=[ck_{i,1},...,ck_{i,j},$...] from the cultural knowledge space, we design an inspection prompt to elicit cultural knowledge by computing a forward pass. The inspection prompt ends with a placeholder token ‘x’, where we perform the patching as in the Patchscope Ghandeharioun et al. (2024). At the $l$-th layer, we replace the hidden representation at the placeholder token position with $x_{*}^{l}$. The inspection prompt can be found in Appendix B.
Step 3. Filtering
We empirically observe that an LLM tends to generate knowledge that is not culture-specific with our inspection prompt when the patched representation lacks the cultural knowledge. Since our method aims at eliciting any cultural knowledge available within inner representations, we devise a filtering method rather than identifying the most relevant model component.
To filter out the knowledge that is unrelated to cultural knowledge, we calculate the semantic similarity as an activation score between the input text $T$ and the generated cultural knowledge $ck_{i,j}$. We take a DeBERTA natural language inference (NLI) model to obtain the hidden representation. [^2] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 In Eq. 3, we compute the representative representation of the input text $T$ using the final hidden states from the NLI model. The activation score $s_{i,j}$ of $ck_{i,j}$ is calculated by their cosine similarity.
$g_{t}^{*}=\frac{1}{T}\sum_{t=0}^{T}g_{t}$ (3)
We keep $ck_{i,j}$ when its $s_{i,j}$ is higher than the threshold, which is set to 0.3.
4.2 Cultural Flattening Score
LLMs trained on imbalanced cultural resources are likely to represent less-documented cultures through the cultural knowledge of more dominant ones, potentially leading to overgeneralization. To quantify this phenomenon, we introduce a Cultural Flattening score (CF score) that measures the extent to which one country’s learned representation has been homogenized to resemble another’s.
A CF score is asymmetric and calculated for a pair of countries, target country $y_{t}$ and source country $y_{s}$. To calculate the CF score, we first compute a cultural knowledge signature for each country. A cultural knowledge signature is a collection of cultural knowledge decoded from Culturescope (§4.1). Let $\mathcal{CK}$ denote the set of cultural knowledge and $\mathcal{Y}$ the set of all countries. For a country $y\in\mathcal{Y}$, let $S_{y}(ck)$ represent the set of activation scores obtained by Eq. 3 for knowledge $ck\in\mathcal{CK}_{y}$ across all instances from country $y$. We define the unnormalized knowledge signature for country $y$ as:
$\tilde{\sigma}_{y}(ck)=\begin{cases}\bar{s}_{y,ck}\cdot\log(1+|S_{y}(ck)|)&\text{if }S_{y}(ck)\neq\emptyset\\ 0&\text{otherwise}\end{cases}$ (4)
where $\bar{s}_{y,ck}=\frac{1}{|S_{y}(ck)|}\sum_{s\in S_{y}(ck)}s$ is the average activation score for knowledge $ck$ in country $y$, and $|S_{y}(ck)|$ denotes the frequency of knowledge $ck$ appearing in country $y$’s data.
The normalized cultural knowledge signature is then computed as:
$\sigma_{y}(ck)=\frac{\tilde{\sigma}_{y}(ck)}{\sum_{ck^{\prime}\in\mathcal{CK}}\tilde{\sigma}_{y}(ck^{\prime})}$ (5)
We define the set of universal concepts as those appearing in every country: $\mathcal{U}=\{ck\in\mathcal{CK}:\forall y\in\mathcal{Y},S_{y}(ck)\neq\emptyset\}$. The CF score from target country $y_{t}$ to source country $y_{s}$ is:
$F(y_{t}\rightarrow y_{s})=\sum_{ck\in\mathcal{CK}\setminus\mathcal{U}}\sigma_{y_{t}}(ck)\cdot\mathbb{1}[S_{y_{s}}(ck)\neq\emptyset]$ (6)
where $\mathbb{1}[\cdot]$ is the indicator function. This formulation ensures that $F(y_{t}\rightarrow y_{s})\in[0,1]$, with higher values indicating that a larger fraction of $y_{t}$’s culturally distinctive knowledge is shared with $y_{s}$.
This asymmetric nature of scores directs cultural influence: a high score $F(y_{t}\rightarrow y_{s})$ indicates that country $y_{t}$’s representation substantially overlaps with country $y_{s}$’s, suggesting that the model may have learned to represent $y_{t}$ through the $y_{s}$’s cultural knowledge.
5 Tracing LLMs’ Internal Mechanisms for Cultural Knowledge Usage
Leveraging MCQs with hard negatives (§3.2.3), we propose to investigate how LLMs internalize Western-dominance bias and cultural flattening via the attention map Yuksekgonul et al. (2024). This setup facilitates an analysis of whether cultural biases are also reflected in the attention mechanism, as in extrinsic evaluation, tracing the internal mechanisms for the emergence of cultural biases.
5.1 Attention Contribution Score
To examine how LLMs internally process Western-dominance bias and cultural flattening, we analyze how attention patterns are directed toward each option $c$ in the input $T$ using the MCQs with hard negatives which contains option list simulating the bias (§3.2.3). Following Yuksekgonul et al. (2024)’s work that highlights the final input token as a meaningful anchor point for attention analysis, we track attention from this final input token $t_{s}$ to tokens that correspond to options in an option list $C$.
For a tokenized input text $T$ that contains a culture-specific question $q$ and a curated option list, we compute the attention contribution $a_{t_{c},t_{s}}$ from the final input token $t_{s}$ to the token $t_{c}$ corresponding to an option $c$, using the following procedure. The operation involves four projection matrices $W^{L}_{Q},W^{L}_{K},W^{L}_{V},W^{L}_{O}\in\mathbb{R}^{d\times d}$ that correspond to the ‘query’, ‘key’, ‘value’, and ‘output’ projections in the attention block of the layer. Each of these is split into multiple heads, where $W^{l,h}_{Q},W^{l,h}_{K},W^{l,h}_{V}\in\mathbb{R}^{d\times d_{h}}$ and $W^{l,h}_{O}\in\mathbb{R}^{d_{h}\times d}$ denote the matrices for each head $h$. $H$ is the total number of heads, $d_{h}$ is the dimensionality for each head. Embeddings are split into equal parts such that $d_{H}=\frac{d}{H}$. The attention weight matrix $A^{l,h}$, calculated as in Eq. 7, is comprised of the attention weight values computed by the attention head $h$ at layer $l$ over an input $T$ containing $S$ number of tokens, where Softmax is taken row-wise. The layer-wise attention contribution score, $a_{t_{c},t_{s}}^{l}$, is defined as in Eq. 8, where $A^{l,h}_{t_{c},t_{s}}$ is the attention weight of the token $t_{c}$ and $t_{s}$ from the matrix $A^{l,h}$.
$A^{l,h}=\mathrm{Softmax}(\frac{(X^{l-1}W^{l,h}_{Q})(X^{l-1}W^{l,h}_{K})^{T}}{\sqrt{d_{h}/H}})$ (7)
$a_{t_{c},t_{s}}^{l}=\sum_{h=1}^{H}A^{l,h}_{t_{c},t_{s}}(x^{l-1}_{t_{c}}W^{l,h}_{V})W^{l,h}_{O}$ (8)
The final attention contribution score, $a_{t_{c}}$, is computed by averaging the layer-wise attention contribution scores $a_{t_{c},t_{s}}^{l}$ across all layers in the LLM, where $t_{s}$ is the final input token.
6 Experimental Results
BLEnD
Llama-3.1 aya-expanse Qwen2.5
Baseline 0.4848 0.4683 0.4626
Cultural Prompting 0.4699 0.4638 0.4664
CANDLE 0.4007 0.2692 0.3473
Culturescope 0.5462 0.4928 0.5059
CAMeL-2
Llama-3.1 aya-expanse Qwen2.5
Baseline 0.7126 0.7019 0.6675
Cultural Prompting 0.6799 0.7282 0.7148
CANDLE 0.5640 0.6955 0.6473
Culturescope 0.6519 0.7169 0.6659
Table 1: We present the QA performances of three different LLMs with different inputs on BLEnD Myung et al. (2024) and CAMeL-2 Naous and Xu (2025). We highlight the best performing method in bold, and the second-best in italics.
6.1 Open-Ended QA Performances
To ensure its relevance to the input, we prepend the comma-separated cultural knowledge $CK$ to the input text $T$. The input example can be found in Appendix B. Given the augmented input $\mathcal{CK}_{i};T_{i}]$, an LLM is asked to generate an output $O$. To evaluate an LLM’s answer $O$, we perform exact-match evaluation, whether $O$ contains a gold answer $a$. We compare our method to three different input schemes, which are designed to enhance LLMs’ cultural understanding ability. Table 1 shows the accuracy from different input schemes.
Baseline is an input without an explicit instruction and additional cultural knowledge. Cultural Prompting Li et al. (2024); Cheng et al. (2023) is an input with an additional instruction (e.g. “The following question is about country yy”) designed to guide an LLM with an explicit country name. CANDLE Nguyen et al. (2023b) is a comprehensive cultural commonsense knowledge base consisting of triples (country, topic, assertion) spanning 196 countries. It provides cultural concepts derived from assertions, which we compare to the cultural knowledge decoded from Culturescope.
Since topics in CANDLE and those in the datasets are not aligned, we sample 20 concepts for each country. Due to the sampling, we perform the inference with three different random seeds and report the average accuracy. Experimental results show that inputs augmented with Culturescope perform better than those with CANDLE concepts. The accuracy with Culturescope is best on BLEnD across all models, on par on CAMeL-2 dataset with the baseline. This confirms that Culturescope reveals a valid cultural knowledge space, which is highly relevant to the given input.
Figure 3: We present the results from the CF score on BLEnD with Llama-3.1. A connection from $y_{t}$ to $y_{s}$ means that the fraction of $y_{t}$’s distinctive knowledge results from the cultural flattening with $y_{s}$.
6.2 Results with CF Score
Figure 3 illustrates the country pairs that exhibit cultural flattening, as measured by our CF score. To highlight only the most significant connections, we apply a threshold to exclude pairs with CF scores lower than the average of all country pairs. Notably, Assam and Indonesia are omitted from the visualization due to their relatively low scores. Overall, the results suggest that countries within the same regional groups tend to share a common cultural knowledge space with bidirectional connections. Additionally, Iran, the United States, and China appear prominently as source countries ($y_{s}$), indicating their broader influence on the connected target countries ($y_{t}$). Further results using CAMeL-2 and other models are available in Appendix C.
BLEnD-Resource BLEnD-Region
Acc % Biased % Others Refusal Acc % Biased % Others Refusal
Llama-3.1 avg. 0.43 0.19 0.18 0.02 0.43 0.20 0.18 0.02
mid 0.44 0.19 0.18 0.02 0.45 0.21 0.16 0.02
low 0.39 0.21 0.19 0.02 0.37 0.19 0.21 0.03
aya-expanse avg. 0.38 0.18 0.16 0.11 0.35 0.19 0.17 0.13
mid 0.40 0.17 0.16 0.11 0.37 0.20 0.15 0.13
low 0.31 0.20 0.19 0.12 0.30 0.19 0.19 0.13
Qwen2.5 avg. 0.44 0.20 0.16 0.05 0.44 0.20 0.16 0.05
mid 0.44 0.19 0.17 0.04 0.47 0.19 0.15 0.04
low 0.40 0.21 0.17 0.06 0.36 0.21 0.18 0.06
Table 2: Model outputs result from Llama-3.1, aya-expanse, and Qwen2.5 on the BLEnD-Resource and BLEnD-Region dataset (§3.2.3), evaluated using four metrics. Results are aggregated by question’s target country type: all averaged (avg.), mid-resource (mid), and low-resource (low).
6.3 Performances on MCQ with Hard Negatives
Table 2 shows the model’s final output results from Llama-3.1, aya-expanse, and Qwen2.5 on the BLEnD-Resource and BLEnD-Region (§3.2.3). We aggregate the results by question’s target country type: all averaged (avg.), mid-resource (mid), and low-resource (low). This reflects LLM’s output-level preference for hard negative options over other options, which serves to reveal how internal biases within the LLM can affect its final prediction. We present the percentage of instances where the model chooses a gold answer, denoted as Accuracy (Acc). The metric labeled as “% Biased” indicates the proportion of instances where the model chooses a hard negative option, which represents a targeted bias. “% Others” represents the proportion of instances where the model chooses one of the remaining random options. Since there are two random options present in the option list, we divide the proportion of choosing random options by two for a fair comparison. We also report the proportion of instances as ‘Refusal’ where LLMs avoid answering.
In most cases, LLMs prefer a hard negative option to other options when they are generating wrong answers with a higher % Biased than % Others. By breaking down the results by resource levels Joshi et al. (2020), we observe that the accuracy decreases for low-resource target questions compared to the average. We also find that in BLEnD-resource, the proportion for selecting a hard negative option from high-resource cultures increases, suggesting the presence of Western-dominance bias. However, this trend does not necessarily extend to BLEnD-region, where the selection rate of hard negatives does not show a similar increase. This indicates that low-resource cultures may be less susceptible to cultural flattening, likely due to the limited cultural knowledge available in the model.
(a) BLEnD-Resource
(b) BLEnD-Region
Figure 4: We present a heatmap visualization of attention contribution scores (z-score normalized) for the Llama-3.1 model on incorrect predictions. Groups on x-axis represent the group of the country which the gold answer is from. Groups on y-axis represent the group of the country which the chosen option is from. For example, in Figure 4 (a), ‘Low (x-axis)’-‘High (y-axis)’ pair with the score of 0.44 shows the averaged attention contribution score on option tokens from ‘High’ resource group when the gold answer is from ‘Low’ resource group. In BLEnD-Resource, we observe high attention contribution scores to tokens from high-resource cultures. BLEnD-Region also shows similar trends, higher scores to Europe and North America, regional group consisting of high-resource countries.
6.4 Attention Contribution Score Analysis
Figure 4 presents the average of attention contribution scores on option token positions, assigned by Llama-3.1 when the model makes incorrect predictions. We separate the analysis between correct predictions and incorrect predictions, as we are particularly interested in LLMs’ internal patterns when they are making biased predictions. Details for the aggregation method and results from aya-expanse and Qwen2.5 are shown in Appendix D.
In both BLEnD-Resource and BLEnD-Region, groups on x-axis represent groups of the question’s target country—resource level (High, Mid, Low) in BLEnD-Resource and region (South Asia [S-AS], East Asia [E-AS], West Asia [W-AS], Europe [EUR], America [AME], Africa [AFR]) in BLEnD-Region — while groups on y-axis indicate groups of the country which the chosen option is from. In Figure 4 (a), visualizing results of BLEnD-resource, we find that attention contributions from the last input token to incorrect high-resource country options are higher than incorrect mid- and low-resource countries options, especially for low-resource target questions. Figure 4 (b), visualizing results of BLEnD-region, shows similar trends to Figure 4 (a). Llama-3.1 allocates higher attention contributions to European and North American countries options compared to other region groups options. This trend is also shown in other two models as in Figure 7.
The analysis of attention contributions demonstrates that high-resource bias and Western-dominance bias is highly internalized within LLMs’ representations. This contradicts the evaluation on LLMs’ performance in Table 2, where LLMs are equally exhibiting both biases. This further aligns with findings from societal biases where they find intrinsic bias metrics and extrinsic bias metrics do not always correlate Goldfarb-Tarrant et al. (2021).
7 Discussion
7.1 Cultural Flattening within LLMs’ Inner Representations
With our Culturescope method, we can now probe the cultural knowledge encoded within the internal representations of LLMs. As shown in Figure 3, which visualizes the cultural flattening direction between cultures, we find unidirectional connections that have Iran and the United States as $y_{s}$. This unidirectional connection implies that models may have learned to represent less documented cultures, such as Ethiopia and Algeria, through those high-resource cultures. Our experiments using hard negative options align with previous works, which find that LLMs sometimes respond with answers aligned with culturally similar or geographically proximate regions Cao et al. (2023); Tao et al. (2024). We further attribute the models’ tendency to favor culturally adjacent answers to the unidirectional connections found by the proposed Culturescope. These findings underscore the need for methods that can disentangle culturally entangled representations, particularly among similar cultures, to enhance the accuracy and cultural appropriateness of LLM outputs.
7.2 LLMs’ Performance on Low-Resource Cultures
Previous studies Azime et al. (2025); Myung et al. (2024); Li et al. (2024) have shown that LLMs often struggle to utilize knowledge relevant to low-resource cultures. Our MCQ results with hard-negative options (Table 2; §6.3) are consistent with these findings, reinforcing the narrative that LLMs underperform on cultural tasks involving low-resource regions. Additionally, the relatively low cultural flattening (CF) scores for low-resource cultures (§6.2) suggest that these models are less prone to generate culturally flattened outputs due to the limited cultural knowledge encoded in their parameters. This is further supported by the reduced ratio of biased answers with region-type hard negatives for low-resource cultures, as shown in Table 2. These findings indicate that LLMs exhibit weaker cultural biases toward low-resource cultures, not because of improved fairness, but due to a lack of cultural representation. Consequently, improving performance on low-resource cultures may require a different approach — one that prioritizes knowledge acquisition over bias mitigation.
8 Conclusion
In this work, we investigate the complex and often biased ways LLMs process cultural knowledge. We introduce Culturescope, a method that leverages activation patching to probe the internal mechanisms of LLMs, allowing us to analyse the cultural knowledge encoded within their layers. In addition, we quantify the phenomenon of ‘cultural flattening’, where LLMs represent less-documented cultures through the concepts of more dominant or geographically close ones, thereby erasing cultural nuances. We conduct our research across three distinct models using culturally-grounded QA datasets, moving beyond isolated extrinsic evaluations to examine the interaction between cultural knowledge within the models’ parametric space. Our analysis reveals that models over-attend Western-centric tokens internally, indicating internalized Western-dominance bias. The results for low-resource cultures suggest that they may not be as susceptible to cultural biases as other cultures. Our findings suggest that future work should develop a tailored approach that considers the impact of bias and resource levels to improve LLMs’ cultural understanding.
Limitations
While our study provides new insights into cultural knowledge in LLMs, there exists limitations. Although a growing number of benchmarks aim to evaluate cultural knowledge in LLMs, few are suitable for our evaluation setup. To meaningfully compare the probed cultural knowledge across cultures, the datasets must maintain a consistent QA format across different cultures. In addition, to ensure that the evaluation captures how LLMs represent culture on a global scale, it is essential to include a geographically diverse set of cultures. However, we emphasize that our proposed method is model- and task-agnostic, and can be applied to any dataset that meets these requirements.
Due to computational constraints, we are reporting results with  8B models, unable to conduct experiments on larger-scale models. In addition, our analysis does not cover all countries globally. This limitation is not due to methodological oversight but rather the lack of publicly available datasets that include culturally grounded QA data for many regions. As such, our findings are necessarily constrained by the scope of existing resources. Finally, our MCQ with hard negatives (§3.2.3) involves a degree of random sampling. While this introduces some variability, we consider it a reasonable trade-off given the prohibitive cost of exhaustively evaluating all possible negative combinations. We mitigate this by ensuring consistency across runs and focusing on aggregate trends rather than individual instances.
Acknowledgements
[…]
Appendix A Dataset Details
Dataset Task Number of Questions Domain List of Countries
BLEnD
Myung et al. (2024)
Cultural
Commonsense QA
5726
Education, Food, Holidays/Celebration/Leisure,
Sport, Work life, Family
Africa: Algeria, Ethiopia, Northern Nigeria
Europe: Spain, United Kingdom, Greece
North America: United States, Mexico
East Asia: China, South Korea
South Asia: Indonesia, Assam
West Asia: Iran, Azerbaijan
CAMeL-2
Naous and Xu (2025)
Extractive QA 1862 Locations, Beverage, Food, Sports
Africa: Morocco, Algeria
Europe: Spain, United Kingdom, Greece
North America: United States, Mexico
East Asia: China, Japan
South Asia: Indonesia, India
West Asia: Iran, Syria, Egypt
Table 4: Details for the datasets
In our experiments, we utilize BLEnD Myung et al. (2024) and CAMeL-2 Naous and Xu (2025). We provide their brief data statistics and characteristics in Table 4.
A.1 Country Groups
As mentioned in Section 3.2, our work conducts experiments that focus on 14 countries classified in two dimensions. We compare three groups based on the level of language resource Joshi et al. (2020) and six groups based on the continental region. Table 3 shows the country entities that correspond to each group.
Appendix B Example of Prompts
In Figure 5, we present the prompt templates we use for each method to obtain Table 1.
Appendix C CF Score Results
In Figure 6, we show the CF score results with Llama-3.1, aya-expanse, Qwen2.5 on BLEnD and CAMeL-2. As mentioned in §6.2, we exclude the countries with CF scores lower than the average CF score across all countries.
(a) BLEnD, aya-expanse
(b) BLEnD, Qwen2.5
(c) CAMeL-2, aya-expanse
(d) CAMeL-2, Qwen2.5
(e) CAMeL-2, Llama-3.1
Figure 6: CF score results
Appendix D Attention Contributions
In cases where an option consists of multiple tokens, we follow the approach of Yuksekgonul et al. (2024), taking the maximum attention contribution score among the component tokens. When the attention scores are averaged by samples to examine its general patterns, simply averaging can be sensitive to extreme values or samples in which attention scores are globally biased toward higher or lower magnitudes. This variability can potentially reduce the generalizability of the results. To address this, we applied normalization (z-score normalization) per sample to the attention contribution scores, such that the scores within each sample have a mean of 0 and a standard deviation of 1.
(a) BLEnD-Resource - Qwen2.5-7B-Instruct
(b) BLEnD-Region - Qwen2.5-7B-Instruct
(c) BLEnD-Resource - Aya Expanse 8B
(d) BLEnD-Region - Aya Expanse 8B
Figure 7: Heatmap visualization of average attention contribution scores (z-score normalized) on incorrect predictions.

### v2

@@
 Abstract
-The growing deployment of large language models (LLMs) across diverse cultural contexts necessitates a better understanding of how the overgeneralization of less documented cultures within LLMs’ representations impacts their cultural understanding.
-Prior work only performs extrinsic evaluation of LLMs’ cultural competence, without accounting for how LLMs’ internal mechanisms lead to cultural (mis)representation.
-To bridge this gap, we propose Culturescope, the first mechanistic interpretability-based method that probes the internal representations of LLMs to elicit the underlying cultural knowledge space.
-CultureScope utilizes a patching method to extract the cultural knowledge.
-We introduce a cultural flattening score as a measure of the intrinsic cultural biases.
-Additionally, we study how LLMs internalize Western-dominance bias and cultural flattening, which allows us to trace how cultural biases emerge within LLMs.
-Our experimental results reveal that LLMs encode Western-dominance bias and cultural flattening in their cultural knowledge space.
-We find that low-resource cultures are less susceptible to cultural biases, likely due to their limited training resources.
+The growing deployment of large language models (LLMs) across diverse cultural contexts necessitates a deeper understanding of LLMs’ representations of different cultures.
+Prior work has focused on evaluating the cultural awareness of LLMs by only examining the text they generate.
+This approach overlooks the internal sources of cultural misrepresentation within the models themselves.
+To bridge this gap, we propose Culturescope, the first mechanistic interpretability-based method that probes the internal representations of different cultural knowledge in LLMs.
+We also introduce a cultural flattening score as a measure of the intrinsic cultural biases of the decoded knowledge from Culturescope.
+Additionally, we study how LLMs internalize cultural biases, which allows us to trace how cultural biases such as Western-dominance bias and cultural flattening emerge within LLMs.
+We find that low-resource cultures are less susceptible to cultural biases, likely due to the model’s limited parametric knowledge.
 Our work provides a foundation for future research on mitigating cultural biases and enhancing LLMs’ cultural understanding.
-Our codes and data used for experiments are publicly available [^1] https://github.com/copenlu/CultureScope.
 1 Introduction
-Large language models (LLMs) are increasingly being used in culturally diverse contexts, where understanding and responding appropriately to each culture is essential (Pandya and Holia, 2023; Salemi et al., 2023; Liu et al., 2025).
-However, the cultural knowledge acquired by LLMs is largely shaped by the data they are trained on, which is predominantly Western-centric (Santurkar et al., 2023).
-This results in severe geographic imbalances, where some regions receive disproportionate attention while others remain marginalized (Nguyen et al., 2023a).
-While LLMs do acquire some cross-cultural patterns during training (Hershcovich et al., 2022; Arora et al., 2023), this learning often results in overgeneralization.
-For example, when asked about a popular leisure activity for retired men in Azerbaijan, GPT-4 responded with chess—a plausible answer, but one that reflects general stereotypes of post-Soviet or Eastern European regions rather than Azerbaijan specifically (Myung et al., 2024).
-Figure 1: Given the question about the popular domestic vacation spot in Greece, Culturescope first generates an answer to cultural Question at the Inference stage.
-Then, it reads the hidden representation from the Inference stage and elicits the cultural knowledge used for ‘Zakynthos’ at the Scoping-in stage.
+Social scientists have long observed that globalization often comes at a cost to cultural diversity.
+This phenomenon, known as ‘Cultural Homogenization,’ occurs when dominant cultures standardize and overgeneralize others, effectively erasing the distinctiveness between cultures Pieterse (1996); Amin (2024).
+Unfortunately, this imbalance is also reflected in the pre-training data of LLMs, which is predominantly Western-centric (Santurkar et al., 2023).
+While LLMs do acquire some cross-cultural knowledge during training (Hershcovich et al., 2022; Arora et al., 2023), the imbalance of the sources of cultural knowledge often leads to cultural bias.
+This results in cultures being either marginalized through underrepresentation or ‘flattened’, wherein their unique traits are conflated with those of dominant cultures.
+(Nguyen et al., 2023a).
+For example, when asked about a popular leisure activity for retired men in Azerbaijan, GPT-4 responded with chess, which is a plausible answer, but one that reflects broad post-Soviet stereotypes rather than culturally specific knowledge about Azerbaijan (Myung et al., 2024).
+Figure 1: Given the question (about the popular domestic vacation spot in Greece), Culturescope first generates an answer to the cultural question at the Inference stage.
+Then, it reads the hidden representation from the Inference stage and elicits the cultural knowledge used for the answer (‘Zakynthos’) at the Scoping-in stage.
 We finalize a list of cultural knowledge after the Filtering stage.
 Culturescope unveils the internal mechanism of LLMs that cannot be revealed through the Inference stage alone.
-However, these patterns of overgeneralization directly perpetuate cultural biases such as cultural flattening.
-Measuring such overgeneralizations through model outputs alone proves insufficient since generated outputs mask the underlying mechanisms that lead to cultural (mis)representation.
-To facilitate the examination of the underlying mechanisms, we propose to study LLMs’ cultural understanding with mechanistic interpretability (MI) techniques.
+These patterns of overgeneralization contribute to what we term ‘cultural flattening’: the model’s tendency to conflate distinct cultural characteristics into dominant cultures.
+While evaluating model outputs, which corresponds to the extrinsic evaluation, is vital for detecting the presence of such overgeneralization, this approach alone cannot reveal the underlying mechanisms that produce it.
+It shows us that a model is biased, but not how this bias is encoded in its parameters.
+To facilitate the examination of the underlying mechanisms, we propose to study LLMs’ cultural biases with mechanistic interpretability (MI) techniques.
 MI techniques provide us with methods that can directly examine how cultural biases discovered by the extrinsic evaluation Santurkar et al.
@@
 We are the first to propose an approach for intrinsic cultural bias evaluation.
-In this work, we introduce Culturescope, a method to probe internal representations and surface the cultural knowledge activated during cultural understanding tasks (§4.1).
-By surfacing this knowledge, we reveal the internal cultural knowledge space from which model outputs are generated.
+In this work, we introduce Culturescope, a method to probe internal representations and surface the cultural biases in play (Sec.
+4.1).
 Figure 1 illustrates an overview of Culturescope.
-To examine the intrinsic ‘cultural flattening’ embedded in the parameter space, we introduce a cultural flattening (CF) score, which quantifies the degree of intersection between cultural knowledge decoded by Culturescope (§4.2).
+To examine the intrinsic cultural flattening embedded in the parameter space, we introduce a cultural flattening (CF) score, which quantifies the degree of intersection between cultural knowledge decoded by Culturescope (Sec.
+4.2).
 We implement our framework on two cultural understanding tasks, cultural commonsense Question Answering (QA) and extractive QA, across three different LLMs.
-We further challenge the model’s cultural understanding by creating multiple-choice questions (MCQs) with hard negatives (§3.2.3).
-Culturally nuanced answers from high-resource cultures or geographically proximate countries are selected as hard negative options to simulate the Western-dominance bias and the cultural flattening.
-This setup refrains LLMs to leverage surface-level elimination strategies based on the overgeneralization Khan et al.
+We further challenge the model’s cultural understanding by creating multiple-choice questions (MCQs) with hard negatives (Sec.
+3.2.3).
+Culturally nuanced answers from high-resource cultures or geographically proximate countries are selected as hard negative options to simulate the cultural biases.
+This setup prevents LLMs from relying on surface-level elimination strategies based on the overgeneralization Khan et al.
 (2025).
 Analyzing selected options by LLMs with the attention map method Yuksekgonul et al.
-(2024) allows us to examine whether extrinsic and intrinsic cultural biases align, by revealing which options the model internally attends to (§5.1).
-High CF scores of Western and high-resource cultures show that LLMs encode Western-dominance bias and cultural flattening in their cultural knowledge space (§6.2).
-Model performances on MCQs with hard negatives demonstrate that low-resource cultures are less likely to be affected by the biases, likely due to their limited training resource (§6.3).
-This implies that LLMs struggle with low-resource cultures due to the lack of parametric knowledge.
-In §6.4, we find that LLMs over-attend to tokens from Western and high-resource cultures.
-This suggests Western-dominance bias is more internalized than the cultural flattening.
-Our work provides useful signal to future research on mitigating the internalized patterns for biases to build a better culturally aligned model.
+(2024) allows us to examine whether extrinsic and intrinsic cultural biases align, by revealing which options the model internally attends to (Sec.
+5.1).
+Our experiments reveal that LLMs internalize a Western-dominance bias and cultural flattening in their cultural knowledge space, as indicated by asymmetric CF scores between cultures (Sec.
+6.1.1) and attention contribution scores (Sec.
+6.1.2).
+This finding is further supported by extrinsic evaluations using MCQs with hard negatives, where LLMs tend to select culturally biased options when answering incorrectly (Sec.
+6.2.1).
+Interestingly, this susceptibility to cultural bias appears weaker for low-resource cultures, likely due to their limited representation in the model’s parametric knowledge.
+Our findings offer valuable insights for future research aimed at mitigating internalized cultural biases and developing culturally aligned LLMs.
 2 Related Work
@@
 (2025); Wang et al.
-(2024); Cao et al.
+(2024a); Cao et al.
 (2024) construct culturally localized evaluation datasets that span domains such as cuisine, proverbs, news, and reasoning.
@@
 (2024).
-Despite growing interest in the cultural capabilities of LLMs, no prior work has explored cultural understanding through the lens of MI.
+Despite growing interest in the cultural capabilities of LLMs, no prior work has explored cultural biases through the lens of MI.
 Our study fills this gap by applying MI techniques to probe the internal representation of cultural knowledge in LLMs, offering new insights into how cultural understanding is encoded and organized within the model.
@@
 3.1 Preliminaries
-A dataset $D=[(q_{1},C_{1},a_{1}),...,(q_{N},C_{N},a_{N})]$ consists of $N$ tuple instances containing: a question $q$, an option list $C=[c_{0},c_{1},c_{2},a]$ containing three options for MCQ, one gold answer $a$ and a country of interest $y$.
-For MCQ, an LLM is given $C$ and $q$ to generate an output $O=[o_{0},...,o_{P}]$ consisting of $P$ tokens.
-For open-ended QA, an LLM is only given $q$ to generate an output $O$.
-To generate an answer, an LLM converts a tokenized input text $T=[t_{0},...,t_{S}]$ containing $S$ number of tokens into $d$-dimenstional vectors using the embedding matrix $E\in\mathbb{R}^{|\mathbb{V}\times d|}$.
+For each culture of interest $y$, a dataset $D=\{(q_{0},C_{0}),...,(q_{N-1},C_{N-1})\}$ consists of $N$ tuple instances that contain a question $q$ and an option list $C=[c_{0},c_{1},c_{2},a]$ of one gold answer $a$ and three other options for MCQ.
+For MCQ, an LLM is given $C$ and $q$ to generate an output $O=[o_{0},...,o_{P-1}]$ consisting of $P$ tokens.
+For open-ended QA, an LLM is only given $q$ to generate the output $O$.
+To generate an answer, an LLM converts a tokenized input text $T=[t_{0},...,t_{S-1}]$ containing $S$ number of tokens into $d$-dimenstional vectors using the embedding matrix $E\in\mathbb{R}^{|\mathbb{V}\times d|}$, where $\mathbb{V}$ denotes the number of vocabularies.
 Then, the vectors are processed through $L$ layers, each containing a multi-head self-attention (MHSA) layer and an MLP layer.
@@
 BLEnD Myung et al.
-(2024) is a hand-crafted benchmark designed to evaluate LLMs’ everyday knowledge across diverse cultures in English.
-It comprises 500 short-answer question-answer pairs for each country, where the answers vary depending on the country’s cultural or regional context.
-To reduce computational cost, we exclude North Korea and West Java, resulting in a final selection of 14 countries from the BLEnD dataset.
+(2024) is a hand-crafted multilingual benchmark designed to evaluate LLMs’ everyday knowledge across diverse cultures.
 CAMeL-2 Naous and Xu (2025) is a bilingual benchmark originally constructed to evaluate LLMs’ entity extraction capabilities on Arabic and English entities.
-With CAMeL-2, an LLM is asked to extract an entity from a context collected from Arabic X/Twitter data according to the specified entity type in the input.
-We take the English partition and reduce the dataset to 14 countries to keep a similar country distribution to the BLEnD dataset.
-Dataset details, including domains and countries covered by the datasets, can be found in Appendix A.
+We use English and Spanish questions from BLEnD and use both Arabic and English from CAMeL-2.
+For both datasets, we select 14 cultural groups to keep a similar culture distribution between two datasets.
+Dataset details can be found in Appendix A.
 3.2.2 Grouping of Cultures
@@
 For the resource dimension, we adopt the taxonomy proposed by Joshi et al.
-(2020), which classifies languages into six levels (0: very low-resource to 5: very high-resource).
-For our experiments, we simplify this into three groups: High (Level 5), Mid (Levels 3-4), and Low (Levels 0-2).
-We assign each country a language resource level based on its most widely spoken language provided by Wikipedia.
+(2020), which we simplify into three resource levels.
 For the region dimension, we group countries into six regions based on continents.
-We split Asia into three subregions, which leaves us six regional groups: North America, Europe, Africa, West Asia, South Asia, and East Asia.
-A complete list of countries within each group is provided in Appendix A.1.
+For more details, we refer to Appendix A.1.
 3.2.3 Cultural MCQ with hard negatives
 Khan et al.
-(2025) found that if MCQs lack the adversarial depth to probe genuine cultural understanding, models can exploit surface-level elimination strategies without truly understanding cultural distinctions.
+(2025) found that if MCQs lack the adversarial depth to probe genuine cultural understanding, models can exploit surface-level elimination strategies without understanding cultural distinctions.
 Thus, we propose a cultural MCQ with hard negatives to study how overgeneralization—driven by regional or resource dominance or similarity—affects the downstream task.
@@
 On the other hand, if the answer is ‘Monopoly’, the LLM is generating an answer from a high resource culture.
-We design two types of multiple choice question option lists that incorporate hard negative options: $C_{resource}$ and $C_{region}$, corresponding to BLEnD-resource and BLEnD-region, respectively.
-For $C_{resource}$, given a question $q$ targeting country $y$, we sample one country from each of the three resource levels excluding $y$.
-We obtain these three country’s respective gold answers when substituted into $q$ for $y$, resulting in three hard negative options: $c_{high}$, $c_{mid}$, and $c_{low}$.
-For $C_{region}$, we sample one country from the same geographical region as $y$ (excluding $y$) and extract its corresponding answer to construct a region-based hard negative option, $c_{flatten}$.
-Two additional options are randomly selected from countries in different regions.
+For dataset creation, we design two types of multiple choice question option lists that incorporate hard negative options: $C_{resource}$ and $C_{region}$, corresponding to BLEnD-resource and BLEnD-region, respectively.
+For $C_{resource}$, given a question $q$ targeting culture $y$, we sample one culture from each of the three resource levels excluding $y$.
+We obtain these three cultures’ respective gold answers when substituted into $q$ for $y$, resulting in three hard negative options: $c_{high}$, $c_{mid}$, and $c_{low}$.
+For $C_{region}$, we sample one culture from the same geographical region as $y$ (excluding $y$) and extract its corresponding answer to construct a region-based hard negative option, $c_{sameregion}$.
+Two additional options ($c_{region1}$, $c_{region2}$) are randomly selected from countries in different regions.
+We report descriptive statistics of the number of selected countries to demonstrate the balanced sampling in Appendix A.2.
 All options are shuffled to avoid positional bias.
@@
 3.3 Models
-Application of MI methods requires a full-access to the model weights.
-Due to the requirement, we conduct our experiments with three recent open-sourced LLMs: Meta-Llama-3.1-8B-Instruct (Llama-3.1, Grattafiori et al.
+Application of MI methods requires full access to model weights, hence we conduct our experiments with three recent open-sourced LLMs: Meta-Llama-3.1-8B-Instruct (Llama-3.1, Grattafiori et al.
 (2024)), aya-expanse-8b (aya-expanse, Dang et al.
@@
 Patchscope Ghandeharioun et al.
-(2024) utilizes an LLM itself to generate natural language explanations of its internal representations.
-It consists of two forward passes, one using a source prompt and the other using an inspection prompt, with a patching operation between them.
-An inspection prompt is designed to guide an LLM as a probe to extract specific knowledge encoded in its internal representations, aligned with a predefined objective, such as next-token prediction, or attribute extraction.
+(2024) utilizes an LLM to generate natural language explanations of its internal representations via two forward passes with a patching operation in between.
+An inspection prompt serves as a probe to extract specific knowledge encoded in the model’s internal states, aligned with a predefined objective, such as next-token prediction, or attribute extraction.
 Utilizing an LLM itself with an inspection prompt as a probing mechanism addresses key limitations of prior methods Hernandez et al.
@@
 (2025), which often rely on predefined probing classes or suffer from limited interpretability due to sub-word tokenization.
-In the context of cultural knowledge, these limitations are particularly pronounced.
-It is challenging to exhaustively define all relevant cultural knowledge.
-Additionally, a direct projection onto the vocabulary space via the unembedding matrix becomes difficult to interpret since cultural knowledge is frequently tokenized into multiple tokens Naous and Xu (2025).
+These limitations are particulary pronounced for cultural knowledge, which is difficult to enumerate exhaustively and often spans multiple tokens Naous and Xu (2025).
 To address these challenges, we introduce a Patchscope-based method tailored for probing the cultural knowledge space.
@@
 (2024).
+Culturescope consists of three stages: inference, scoping-in, and filtering.
+Culturescope allows us to move beyond what is observable from model responses alone, overcoming the limitation of extrinsic evaluation.
 4.1 Culturescope
-Culturescope consists of three stages: inference, scoping-in, and filtering.
-Our Culturescope allows us to move beyond what is observable from model responses alone, overcoming the limitation of extrinsic evaluation.
 Step 1. Inference
-An LLM first encodes a tokenized input $T_{i}$ of $i$-th instance and generates an output $O_{i}$, which is an LLM answer to an Open-Ended cultural QA.
-Then, we compute the representative hidden representations used to generate an answer, which are patched onto the inspection prompt during the scoping-in stage.
-As Patchscope does not consider patching with multi-tokens Ghandeharioun et al.
+An LLM first encodes a tokenized input $T_{i}$ of $i$-th instance and generates an output $O_{i}$, which is an LLM answer to an open-ended cultural QA consisting of $P$ number of tokens.
+However, activation patching during the scoping-in stage requires a single-token representation of the $O_{i}$.
+Since Patchscope does not consider patching with multi-tokens Ghandeharioun et al.
 (2024), we adopt Bronzini et al.
-(2024)’s approach originally developed for fact-checking claims to condense an LLM’s cultural answer involving multiple tokens into a single hidden representation.
-Similarly, we compute the representative hidden representation $x_{*}^{l}$ of the $O_{i}$, which consists of multiple tokens, on the layer $l$.
-Specifically, we perform the weighted sum of hidden states as in Eq.
+(2024)’s approach, originally developed for fact-checking claims, to condense an LLM’s cultural answer involving multiple tokens into a single hidden representation.
+Specifically, to compute the representative hidden representation $x_{*}^{l}$, we perform the weighted sum of hidden states as in Eq.
 2 for the layer $l$.
@@
 Other token weights are set to zero.
-We use this $x_{*}^{l}$ in the next stage.
-$x_{*}^{l}=\sum_{p=0}^{P}x_{p}^{l}*w_{p}$ (2)
+The resulting $x_{*}^{l}$ is then patched onto the inspection prompt during the scoping-in stage.
+$x_{*}^{l}=\sum_{p=0}^{P-1}x_{p}^{l}*w_{p}$ (2)
 Step 2. Scoping-in
-At this stage, we further elicit the cultural knowledge encoded in $x_{*}^{l}$ to reveal the cultural knowledge space utilized for the $O_{i}$.
-To elicit a list of cultural knowledge $\mathcal{CK}_{i}=[ck_{i,1},...,ck_{i,j},$...] from the cultural knowledge space, we design an inspection prompt to elicit cultural knowledge by computing a forward pass.
-The inspection prompt ends with a placeholder token ‘x’, where we perform the patching as in the Patchscope Ghandeharioun et al.
+During this stage, we utilize activation patching to elicit the cultural knowledge encoded in $x_{*}^{l}$, revealing the cultural knowledge space utilized for generating $O_{i}$.
+To do this, the LLM performs another inference using an inspection prompt, which generates a comma-separated sequence of cultural knowledge items.
+The inspection prompt ends with a placeholder token ‘x’, where patching is done following Patchscope Ghandeharioun et al.
 (2024).
-At the $l$-th layer, we replace the hidden representation at the placeholder token position with $x_{*}^{l}$.
+At layer $l$, we replace the hidden representation at the placeholder token position with $x_{*}^{l}$.
 The inspection prompt can be found in Appendix B.
+Finally, we split the generated sequence into a list of cultural knowledge $\mathcal{K}_{i}=[k_{i,1},...,k_{i,j},$...].
 Step 3. Filtering
@@
 Since our method aims at eliciting any cultural knowledge available within inner representations, we devise a filtering method rather than identifying the most relevant model component.
-To filter out the knowledge that is unrelated to cultural knowledge, we calculate the semantic similarity as an activation score between the input text $T$ and the generated cultural knowledge $ck_{i,j}$.
-We take a DeBERTA natural language inference (NLI) model to obtain the hidden representation.
-[^2] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 In Eq.
-3, we compute the representative representation of the input text $T$ using the final hidden states from the NLI model.
-The activation score $s_{i,j}$ of $ck_{i,j}$ is calculated by their cosine similarity.
-$g_{t}^{*}=\frac{1}{T}\sum_{t=0}^{T}g_{t}$ (3)
-We keep $ck_{i,j}$ when its $s_{i,j}$ is higher than the threshold, which is set to 0.3.
+To filter out the knowledge that is unrelated to cultural knowledge, we calculate the semantic similarity as an activation score between the input text $T$ and the generated cultural knowledge $k_{i,j}$.
+We take separate encoder-based semantic search models for each language to obtain the hidden representation.
+In Eq.
+3, we compute the representation of the input text $H_{T}=[h_{0},..h_{s},..h_{S-1}]$ using the final hidden states from the semantic search model.
+The activation score $e_{i,j}$ of $k_{i,j}$ is calculated by the cosine similarity.
+$h_{t}^{*}=\frac{1}{S}\sum_{s=0}^{S-1}h_{t}$ (3)
+We keep $k_{i,j}$ when its $e_{i,j}$ is higher than the average of all $e$. Details for implementation is provided in Appendix C.
 4.2 Cultural Flattening Score
-LLMs trained on imbalanced cultural resources are likely to represent less-documented cultures through the cultural knowledge of more dominant ones, potentially leading to overgeneralization.
-To quantify this phenomenon, we introduce a Cultural Flattening score (CF score) that measures the extent to which one country’s learned representation has been homogenized to resemble another’s.
-A CF score is asymmetric and calculated for a pair of countries, target country $y_{t}$ and source country $y_{s}$.
-To calculate the CF score, we first compute a cultural knowledge signature for each country.
-A cultural knowledge signature is a collection of cultural knowledge decoded from Culturescope (§4.1).
-Let $\mathcal{CK}$ denote the set of cultural knowledge and $\mathcal{Y}$ the set of all countries.
-For a country $y\in\mathcal{Y}$, let $S_{y}(ck)$ represent the set of activation scores obtained by Eq.
-3 for knowledge $ck\in\mathcal{CK}_{y}$ across all instances from country $y$.
-We define the unnormalized knowledge signature for country $y$ as:
-$\tilde{\sigma}_{y}(ck)=\begin{cases}\bar{s}_{y,ck}\cdot\log(1+|S_{y}(ck)|)&\text{if }S_{y}(ck)\neq\emptyset\\ 0&\text{otherwise}\end{cases}$ (4)
-where $\bar{s}_{y,ck}=\frac{1}{|S_{y}(ck)|}\sum_{s\in S_{y}(ck)}s$ is the average activation score for knowledge $ck$ in country $y$, and $|S_{y}(ck)|$ denotes the frequency of knowledge $ck$ appearing in country $y$’s data.
-The normalized cultural knowledge signature is then computed as:
-$\sigma_{y}(ck)=\frac{\tilde{\sigma}_{y}(ck)}{\sum_{ck^{\prime}\in\mathcal{CK}}\tilde{\sigma}_{y}(ck^{\prime})}$ (5)
-We define the set of universal concepts as those appearing in every country: $\mathcal{U}=\{ck\in\mathcal{CK}:\forall y\in\mathcal{Y},S_{y}(ck)\neq\emptyset\}$.
-The CF score from target country $y_{t}$ to source country $y_{s}$ is:
-$F(y_{t}\rightarrow y_{s})=\sum_{ck\in\mathcal{CK}\setminus\mathcal{U}}\sigma_{y_{t}}(ck)\cdot\mathbb{1}[S_{y_{s}}(ck)\neq\emptyset]$ (6)
+To quantify cultural flattening, where a culture’s distinctive characteristics are misrepresented by those of a dominant culture, we introduce the Cultural Flattening score (CF score).
+A CF score is asymmetric and is calculated for an ordered pair of cultures, a target culture $y_{t}$ and a source culture $y_{s}$.
+For the CF score, we first compute a chi-square contribution for each cultural knowledge decoded by Culturescope (Sec.
+4.1).
+This chi-square contribution quantifies how strongly a given piece of knowledge is activated for a specific culture relative to an independence assumption across all cultures, thereby serving as a measure of cultural specificity.
+This formulation highlights the impact of culturally distinctive knowledge rather than shared cultural knowledge arising from cultural assimilation.
+For knowledge $k$ from a culture $y\in\mathcal{Y}$, the expected count under the independence assumption, $\mathbb{E}_{k,y}$, is defined as:
+$\mathbb{E}_{k,y}=\frac{N_{y}\cdot\sum_{y}g_{k,y}}{N_{\mathcal{Y}}}$ (4)
+where $g_{k,y}$ is the frequency of $k$ in the culture $y$, and $N$ denotes the number of knowledge generated by $y$ or across all cultures $\mathcal{Y}$.
+The chi-square contribution of cultural knowledge $\mathcal{X}_{k,y}$ is then computed as:
+$\mathcal{X}_{k,y}=\frac{(max(0,g_{k,y}-\mathbb{E}_{k,y}))^{2}}{\mathbb{E}_{k,y}}$ (5)
+We only consider cases where the residuals are positive, in order to remove knowledge that is underrepresented by the culture.
+The higher the $\mathcal{X}_{k,y}$, the more distincitve $k$ is.
+The final CF score from the target culture $y_{t}$ to the source culture $y_{s}$ is:
+$F(y_{t}\rightarrow y_{s})=\sum_{k\in\mathcal{K}}\mathcal{X}_{k,y_{t}}\cdot\mathbb{1}[\mathcal{X}_{k,y_{s}}\neq 0]$ (6)
 where $\mathbb{1}[\cdot]$ is the indicator function.
-This formulation ensures that $F(y_{t}\rightarrow y_{s})\in[0,1]$, with higher values indicating that a larger fraction of $y_{t}$’s culturally distinctive knowledge is shared with $y_{s}$.
-This asymmetric nature of scores directs cultural influence: a high score $F(y_{t}\rightarrow y_{s})$ indicates that country $y_{t}$’s representation substantially overlaps with country $y_{s}$’s, suggesting that the model may have learned to represent $y_{t}$ through the $y_{s}$’s cultural knowledge.
+By summing chi-square contribution scores over overlapping knowledge, the CF score captures the total amount of culturally distinctive knowledge from the target culture that is also present in the source culture.
+Higher values indicate that a large fraction of the target culture’s distinctive concepts are represented in the source culture.
 5 Tracing LLMs’ Internal Mechanisms for Cultural Knowledge Usage
-Leveraging MCQs with hard negatives (§3.2.3), we propose to investigate how LLMs internalize Western-dominance bias and cultural flattening via the attention map Yuksekgonul et al.
+Leveraging MCQs with hard negatives (Sec.
+3.2.3), we propose to investigate how LLMs internalize Western-dominance bias and cultural flattening via the attention map Yuksekgonul et al.
 (2024).
@@
 5.1 Attention Contribution Score
-To examine how LLMs internally process Western-dominance bias and cultural flattening, we analyze how attention patterns are directed toward each option $c$ in the input $T$ using the MCQs with hard negatives which contains option list simulating the bias (§3.2.3).
+To examine how LLMs internally process Western-dominance bias and cultural flattening, we analyze how attention patterns are directed toward each option $c$ in the input $T$ using the MCQs with hard negatives (Sec.
+3.2.3).
 Following Yuksekgonul et al.
-(2024)’s work that highlights the final input token as a meaningful anchor point for attention analysis, we track attention from this final input token $t_{s}$ to tokens that correspond to options in an option list $C$.
-For a tokenized input text $T$ that contains a culture-specific question $q$ and a curated option list, we compute the attention contribution $a_{t_{c},t_{s}}$ from the final input token $t_{s}$ to the token $t_{c}$ corresponding to an option $c$, using the following procedure.
-The operation involves four projection matrices $W^{L}_{Q},W^{L}_{K},W^{L}_{V},W^{L}_{O}\in\mathbb{R}^{d\times d}$ that correspond to the ‘query’, ‘key’, ‘value’, and ‘output’ projections in the attention block of the layer.
-Each of these is split into multiple heads, where $W^{l,h}_{Q},W^{l,h}_{K},W^{l,h}_{V}\in\mathbb{R}^{d\times d_{h}}$ and $W^{l,h}_{O}\in\mathbb{R}^{d_{h}\times d}$ denote the matrices for each head $h$.
-$H$ is the total number of heads, $d_{h}$ is the dimensionality for each head.
-Embeddings are split into equal parts such that $d_{H}=\frac{d}{H}$.
-The attention weight matrix $A^{l,h}$, calculated as in Eq.
-7, is comprised of the attention weight values computed by the attention head $h$ at layer $l$ over an input $T$ containing $S$ number of tokens, where Softmax is taken row-wise.
-The layer-wise attention contribution score, $a_{t_{c},t_{s}}^{l}$, is defined as in Eq.
-8, where $A^{l,h}_{t_{c},t_{s}}$ is the attention weight of the token $t_{c}$ and $t_{s}$ from the matrix $A^{l,h}$.
-$A^{l,h}=\mathrm{Softmax}(\frac{(X^{l-1}W^{l,h}_{Q})(X^{l-1}W^{l,h}_{K})^{T}}{\sqrt{d_{h}/H}})$ (7)
-$a_{t_{c},t_{s}}^{l}=\sum_{h=1}^{H}A^{l,h}_{t_{c},t_{s}}(x^{l-1}_{t_{c}}W^{l,h}_{V})W^{l,h}_{O}$ (8)
-The final attention contribution score, $a_{t_{c}}$, is computed by averaging the layer-wise attention contribution scores $a_{t_{c},t_{s}}^{l}$ across all layers in the LLM, where $t_{s}$ is the final input token.
+(2024)’s work that highlights the final input token as a meaningful anchor point for attention analysis, we track attention from this final input token $t_{s}$ to tokens that correspond to $c$ in an option list $C$.
+To avoid biases from MCQ Wang et al.
+(2024b), we provide options without option letters.
+For a tokenized input text $T$ that contains a culture-specific question $q$ and a curated option list, we compute the attention contribution $a_{t_{c},t_{s}}$ from the final input token $t_{s}$ to the token $t_{c}$ corresponding to an option $c$, using the attention weight of the token $t_{c}$ and $t_{s}$.
+As in Eq.
+7, we obtain a layer-wise attention contribution score $a_{t_{c},t_{s}}^{l}$ on layer $l$, where $ATTN_{t_{c},t_{s}}$ denotes attention operation between the token $t_{c}$ and $t_{s}$ with a hidden representation $x^{l-1}_{t_{c}}$ from the previous layer.
+$a_{t_{c},t_{s}}^{l}=ATTN_{t_{c},t_{s}}(x^{l-1}_{t_{c}})$ (7)
+The final attention contribution score, $a_{t_{c}}$, is computed by averaging the layer-wise scores $a_{t_{c},t_{s}}^{l}$ across all layers in the LLM.
+Figure 3: We present the results from the CF score on BLEnD with Llama-3.1 in English.
+Percentages on the right reflect relative knowledge coverage and are defined as the ratio between the number of cultural knowledge generated for a given culture and the total number of generated concepts across all cultures.
 6 Experimental Results
-BLEnD
+6.1 Intrinsic Cultural Bias Evaluation
+6.1.1 CF Score Results
+We present CF score results using a Sankey diagram, where the flow direction indicates target-to-source relations and flow width is proportional to the CF score.
+Figure 3 shows cultural pairs with high CF scores, where strand width reflects the extent to which culturally distinctive knowledge from a target culture $y_{t}$ is generalized to a source culture $y_{s}$ in the model’s internal representations.
+Despite their large number of generated concepts, the US and the UK exhibit relatively weak incoming connections as source countries, indicating that the CF score successfully downweights overlap driven by broadly shared cultural knowledge.
+The result reveals asymmetric connections, often aligned with geographic or cultural proximity, suggesting uneven generalization of culturally distinctive knowledge (China $\rightarrow$ S.
+Korea, Iran $\rightarrow$ Azerbaijan).
+In addition, cultures with lower concept coverage, such as Assam, Azerbaijan, and Ethiopia, consistently show narrow connections.
+This pattern of asymmetry is also shown in other languages.
+Additional results using CAMeL-2 and other models, covering English, Arabic, and Spanish, are provided in Appendix E.
+(a) BLEnD-Resource
+(b) BLEnD-Region
+Figure 4: We present a heatmap of attention contribution scores (z-score normalized) for incorrect predictions of Llama-3.1.
+The x-axis groups correspond to the culture of the chosen option, while the y-axis groups correspond to the culture of the gold answer.
+For example, in Figure 4 (a), the High-Low cell (0.44) indicates the average attention contribution to option tokens from the High resource group when the gold answer belongs to the Low resource group.
+BLEnD-Resource BLEnD-Region
+Acc % Biased % Others Refusal Acc % Biased % Others Refusal
+Llama-3.1 avg. 0.43 0.19** 0.18 0.02 0.43 0.20** 0.18 0.02
+mid 0.44 0.19 0.18 0.02 0.45 0.21** 0.16 0.02
+low 0.39 0.21 0.19 0.02 0.37 0.19 0.21 0.03
+aya-expanse avg. 0.38 0.18** 0.16 0.11 0.35 0.19** 0.17 0.13
+mid 0.40 0.17 0.16 0.11 0.37 0.20** 0.15 0.13
+low 0.31 0.20 0.19 0.12 0.30 0.19 0.19 0.13
+Qwen2.5 avg. 0.44 0.20** 0.16 0.05 0.44 0.20** 0.16 0.05
+mid 0.44 0.19 0.17 0.04 0.47 0.19** 0.15 0.04
+low 0.40 0.21** 0.17 0.06 0.36 0.21** 0.18 0.06
+Table 1: Model outputs result from Llama-3.1, aya-expanse, and Qwen2.5 on the BLEnD-Resource and BLEnD-Region dataset (Sec.
+3.2.3), evaluated using four metrics.
+Results are aggregated by question’s target culture type: all averaged (avg.), mid-resource (mid), and low-resource (low).
+** denotes statistically significant difference between % Biased and % Others.
+6.1.2 Attention Contribution Score Results
+Figure 4 presents the average of attention contribution scores on option token positions, assigned by Llama-3.1 when the model makes incorrect predictions.
+We separate the analysis between correct predictions and incorrect predictions, as we are particularly interested in LLMs’ internal patterns when they are making biased predictions.
+Details for the aggregation method and results from aya-expanse and Qwen2.5 are shown in Appendix F.
+In BLEnD-Resource and BLEnD-Region, the x-axis groups correspond to the groups of the chosen option, while the y-axis groups represent the target culture of the question—resource level (High, Mid, Low) in BLEnD-Resource and region (South Asia [S-AS], East Asia [E-AS], West Asia [W-AS], Europe [EUR], America [AME], Africa [AFR]) in BLEnD-Region.
+From Figure 4 (a) and (b), attention contributions from the last input token to incorrect options reflect the induced biases, Western-dominance bias in BLEnD-Resource and cultural flattening in BLEnD-Region.
+This trend is also shown in two additional models (Figure 8).
+Student’s t-tests indicate that the average differences in attention scores are statistically significant.
+Details are provided in Appendix D.2.
+6.2 Extrinsic Cultural Bias Evaluation
+6.2.1 Performances on MCQ with Hard Negatives
+Table 1 shows the extrinsic evaluation results from Llama-3.1, aya-expanse, and Qwen2.5 on the BLEnD-Resource and BLEnD-Region (Sec.
+3.2.3).
+We aggregate the results by question’s target culture type: all averaged (avg.), mid-resource (mid), and low-resource (low).
+We present the percentage of instances where the model chooses a gold answer, denoted as Accuracy (Acc).
+The metric labeled as “% Biased” indicates the proportion of instances where the model chooses a hard negative option, which represents a targeted bias.
+“% Others” represents the proportion of instances where the model chooses one of the remaining random options.
+Since there are two random options present in the option list, we divide the proportion of choosing random options by two for a fair comparison.
+We also report the proportion of instances as ‘Refusal’ where LLMs avoid answering.
+We observe that the accuracy decreases for low-resource target questions compared to the average, which is statistically significant (Table 7).
+We further statistically test whether the lower accuracy in low-resource cultures stems from a preference for biased options over others.
+However, while the differences between % Biased and % Others are statistically significant in avg.
+and mid-resource cultures, the differences in low-resource cultures are not statistically significant.
+Details for the statistical tests are described in Appendix D.1.
+7 Evaluation of Culturescope
+Considering that Culturescope elicits cultural knowledge from all layers without pinpointing any specific model component, we design two validation approaches including relevance evaluation and irrelevant patching.
+First, for the relevance evaluation, we compare our method to three different input schemes, which are designed to enhance LLMs’ cultural understanding ability.
+Specifically, For Culturescope, we prepend the comma-separated cultural knowledge $K$ to the input text $T$ and evaluate an LLM’s answer $O$ by exact-match.
+Table 2 shows the accuracy from different input schemes on BLEnD.
+Results on CAMeL-2 are in Table 9.
+Baseline is an input without an explicit instruction and additional cultural knowledge.
+Cultural Prompting Li et al.
+(2024); Cheng et al.
+(2023) is an input with an additional instruction (e.g.
+“The following question is about culture yy”) designed to guide an LLM with an explicit culture name.
+CANDLE Nguyen et al.
+(2023b) is a comprehensive cultural commonsense knowledge base which provides cultural concepts derived from assertions.
+The input example can be found in Appendix B.
+Since topics in CANDLE and those in the datasets are not aligned, we sample 20 concepts for each culture.
+Due to the sampling, we perform the inference with three different random seeds and report the average accuracy.
+Experimental results show that inputs augmented with Culturescope perform better than those with CANDLE concepts.
+The accuracy with Culturescope is best on BLEnD across all models, on par on CAMeL-2 dataset with the baseline.
+This confirms that Culturescope reveals a valid cultural knowledge space, which is highly relevant to the given input.
+For the irrelevant patching test, instead of patching representative hidden representations, we patch randomly sampled Gausian noise vectors and decode knowledge using the same inspection prompt.
+We report the average semantic similarity score of the generated knowledge computed during the filtering stage.
+Higher similarity scores for Culturescope validate that the generated cultural knowledge is not an artifact induced by the experimental setup (Table 2; Table 8).
+In addition, we provide the average number of generated knowledge per layer in Appendix C.
+Relevance evaluation
 Llama-3.1 aya-expanse Qwen2.5
@@
 Culturescope 0.5462 0.4928 0.5059
-CAMeL-2
+Irrelevant patching
 Llama-3.1 aya-expanse Qwen2.5
-Baseline 0.7126 0.7019 0.6675
-Cultural Prompting 0.6799 0.7282 0.7148
-CANDLE 0.5640 0.6955 0.6473
-Culturescope 0.6519 0.7169 0.6659
-Table 1: We present the QA performances of three different LLMs with different inputs on BLEnD Myung et al.
-(2024) and CAMeL-2 Naous and Xu (2025).
-We highlight the best performing method in bold, and the second-best in italics.
-6.1 Open-Ended QA Performances
-To ensure its relevance to the input, we prepend the comma-separated cultural knowledge $CK$ to the input text $T$.
-The input example can be found in Appendix B.
-Given the augmented input $\mathcal{CK}_{i};T_{i}]$, an LLM is asked to generate an output $O$.
-To evaluate an LLM’s answer $O$, we perform exact-match evaluation, whether $O$ contains a gold answer $a$.
-We compare our method to three different input schemes, which are designed to enhance LLMs’ cultural understanding ability.
-Table 1 shows the accuracy from different input schemes.
-Baseline is an input without an explicit instruction and additional cultural knowledge.
-Cultural Prompting Li et al.
-(2024); Cheng et al.
-(2023) is an input with an additional instruction (e.g.
-“The following question is about country yy”) designed to guide an LLM with an explicit country name.
-CANDLE Nguyen et al.
-(2023b) is a comprehensive cultural commonsense knowledge base consisting of triples (country, topic, assertion) spanning 196 countries.
-It provides cultural concepts derived from assertions, which we compare to the cultural knowledge decoded from Culturescope.
-Since topics in CANDLE and those in the datasets are not aligned, we sample 20 concepts for each country.
-Due to the sampling, we perform the inference with three different random seeds and report the average accuracy.
-Experimental results show that inputs augmented with Culturescope perform better than those with CANDLE concepts.
-The accuracy with Culturescope is best on BLEnD across all models, on par on CAMeL-2 dataset with the baseline.
-This confirms that Culturescope reveals a valid cultural knowledge space, which is highly relevant to the given input.
-Figure 3: We present the results from the CF score on BLEnD with Llama-3.1.
-A connection from $y_{t}$ to $y_{s}$ means that the fraction of $y_{t}$’s distinctive knowledge results from the cultural flattening with $y_{s}$.
-6.2 Results with CF Score
-Figure 3 illustrates the country pairs that exhibit cultural flattening, as measured by our CF score.
-To highlight only the most significant connections, we apply a threshold to exclude pairs with CF scores lower than the average of all country pairs.
-Notably, Assam and Indonesia are omitted from the visualization due to their relatively low scores.
-Overall, the results suggest that countries within the same regional groups tend to share a common cultural knowledge space with bidirectional connections.
-Additionally, Iran, the United States, and China appear prominently as source countries ($y_{s}$), indicating their broader influence on the connected target countries ($y_{t}$).
-Further results using CAMeL-2 and other models are available in Appendix C.
-BLEnD-Resource BLEnD-Region
-Acc % Biased % Others Refusal Acc % Biased % Others Refusal
-Llama-3.1 avg. 0.43 0.19 0.18 0.02 0.43 0.20 0.18 0.02
-mid 0.44 0.19 0.18 0.02 0.45 0.21 0.16 0.02
-low 0.39 0.21 0.19 0.02 0.37 0.19 0.21 0.03
-aya-expanse avg. 0.38 0.18 0.16 0.11 0.35 0.19 0.17 0.13
-mid 0.40 0.17 0.16 0.11 0.37 0.20 0.15 0.13
-low 0.31 0.20 0.19 0.12 0.30 0.19 0.19 0.13
-Qwen2.5 avg. 0.44 0.20 0.16 0.05 0.44 0.20 0.16 0.05
-mid 0.44 0.19 0.17 0.04 0.47 0.19 0.15 0.04
-low 0.40 0.21 0.17 0.06 0.36 0.21 0.18 0.06
-Table 2: Model outputs result from Llama-3.1, aya-expanse, and Qwen2.5 on the BLEnD-Resource and BLEnD-Region dataset (§3.2.3), evaluated using four metrics.
-Results are aggregated by question’s target country type: all averaged (avg.), mid-resource (mid), and low-resource (low).
-6.3 Performances on MCQ with Hard Negatives
-Table 2 shows the model’s final output results from Llama-3.1, aya-expanse, and Qwen2.5 on the BLEnD-Resource and BLEnD-Region (§3.2.3).
-We aggregate the results by question’s target country type: all averaged (avg.), mid-resource (mid), and low-resource (low).
-This reflects LLM’s output-level preference for hard negative options over other options, which serves to reveal how internal biases within the LLM can affect its final prediction.
-We present the percentage of instances where the model chooses a gold answer, denoted as Accuracy (Acc).
-The metric labeled as “% Biased” indicates the proportion of instances where the model chooses a hard negative option, which represents a targeted bias.
-“% Others” represents the proportion of instances where the model chooses one of the remaining random options.
-Since there are two random options present in the option list, we divide the proportion of choosing random options by two for a fair comparison.
-We also report the proportion of instances as ‘Refusal’ where LLMs avoid answering.
-In most cases, LLMs prefer a hard negative option to other options when they are generating wrong answers with a higher % Biased than % Others.
-By breaking down the results by resource levels Joshi et al.
-(2020), we observe that the accuracy decreases for low-resource target questions compared to the average.
-We also find that in BLEnD-resource, the proportion for selecting a hard negative option from high-resource cultures increases, suggesting the presence of Western-dominance bias.
-However, this trend does not necessarily extend to BLEnD-region, where the selection rate of hard negatives does not show a similar increase.
-This indicates that low-resource cultures may be less susceptible to cultural flattening, likely due to the limited cultural knowledge available in the model.
-(a) BLEnD-Resource
-(b) BLEnD-Region
-Figure 4: We present a heatmap visualization of attention contribution scores (z-score normalized) for the Llama-3.1 model on incorrect predictions.
-Groups on x-axis represent the group of the country which the gold answer is from.
-Groups on y-axis represent the group of the country which the chosen option is from.
-For example, in Figure 4 (a), ‘Low (x-axis)’-‘High (y-axis)’ pair with the score of 0.44 shows the averaged attention contribution score on option tokens from ‘High’ resource group when the gold answer is from ‘Low’ resource group.
-In BLEnD-Resource, we observe high attention contribution scores to tokens from high-resource cultures.
-BLEnD-Region also shows similar trends, higher scores to Europe and North America, regional group consisting of high-resource countries.
-6.4 Attention Contribution Score Analysis
-Figure 4 presents the average of attention contribution scores on option token positions, assigned by Llama-3.1 when the model makes incorrect predictions.
-We separate the analysis between correct predictions and incorrect predictions, as we are particularly interested in LLMs’ internal patterns when they are making biased predictions.
-Details for the aggregation method and results from aya-expanse and Qwen2.5 are shown in Appendix D.
-In both BLEnD-Resource and BLEnD-Region, groups on x-axis represent groups of the question’s target country—resource level (High, Mid, Low) in BLEnD-Resource and region (South Asia [S-AS], East Asia [E-AS], West Asia [W-AS], Europe [EUR], America [AME], Africa [AFR]) in BLEnD-Region — while groups on y-axis indicate groups of the country which the chosen option is from.
-In Figure 4 (a), visualizing results of BLEnD-resource, we find that attention contributions from the last input token to incorrect high-resource country options are higher than incorrect mid- and low-resource countries options, especially for low-resource target questions.
-Figure 4 (b), visualizing results of BLEnD-region, shows similar trends to Figure 4 (a).
-Llama-3.1 allocates higher attention contributions to European and North American countries options compared to other region groups options.
-This trend is also shown in other two models as in Figure 7.
-The analysis of attention contributions demonstrates that high-resource bias and Western-dominance bias is highly internalized within LLMs’ representations.
-This contradicts the evaluation on LLMs’ performance in Table 2, where LLMs are equally exhibiting both biases.
-This further aligns with findings from societal biases where they find intrinsic bias metrics and extrinsic bias metrics do not always correlate Goldfarb-Tarrant et al.
-(2021).
+English 0.325 $\pm$ 0.01 0.349 $\pm$ 0.1 0.344 $\pm$ 0.11
+Irrelevant 0.149 $\pm$ 0.09 0.127 $\pm$ 0.07 0.141 $\pm$ 0.09
+Table 2: We present the evaluation of Culturescope on BLEnD Myung et al.
+(2024) in English.
+For the relevance evaluation results, we highlight the best-performing method in bold and the second-best in underline.
+For the irrelevant patching results, we report the mean and standard deviation.
 8 Discussion
-7.1 Cultural Flattening within LLMs’ Inner Representations
-With our Culturescope method, we can now probe the cultural knowledge encoded within the internal representations of LLMs.
-As shown in Figure 3, which visualizes the cultural flattening direction between cultures, we find unidirectional connections that have Iran and the United States as $y_{s}$.
-This unidirectional connection implies that models may have learned to represent less documented cultures, such as Ethiopia and Algeria, through those high-resource cultures.
-Our experiments using hard negative options align with previous works, which find that LLMs sometimes respond with answers aligned with culturally similar or geographically proximate regions Cao et al.
-(2023); Tao et al.
-(2024).
-We further attribute the models’ tendency to favor culturally adjacent answers to the unidirectional connections found by the proposed Culturescope.
-These findings underscore the need for methods that can disentangle culturally entangled representations, particularly among similar cultures, to enhance the accuracy and cultural appropriateness of LLM outputs.
-7.2 LLMs’ Performance on Low-Resource Cultures
-Previous studies Azime et al.
-(2025); Myung et al.
-(2024); Li et al.
-(2024) have shown that LLMs often struggle to utilize knowledge relevant to low-resource cultures.
-Our MCQ results with hard-negative options (Table 2; §6.3) are consistent with these findings, reinforcing the narrative that LLMs underperform on cultural tasks involving low-resource regions.
-Additionally, the relatively low cultural flattening (CF) scores for low-resource cultures (§6.2) suggest that these models are less prone to generate culturally flattened outputs due to the limited cultural knowledge encoded in their parameters.
-This is further supported by the reduced ratio of biased answers with region-type hard negatives for low-resource cultures, as shown in Table 2.
-These findings indicate that LLMs exhibit weaker cultural biases toward low-resource cultures, not because of improved fairness, but due to a lack of cultural representation.
-Consequently, improving performance on low-resource cultures may require a different approach — one that prioritizes knowledge acquisition over bias mitigation.
+Our proposed methods for evaluating intrinsic cultural biases in LLMs reveal that models internalize two cultural biases: Wester-dominance bias and cultural flattening.
+As shown in Figure 3, we observe asymmetric connections among geographically or culturally proximate cultures.
+These asymmetries indicate uneven generalization of culturally distinctive knowledge, consistent with potential cultural flattening effects.
+Attention-based analyses further reveal the presence of Western-dominance bias that is not captured by CF scores as the CF score explicitly downweights globally shared knowledge (Section 6.1.2; Figure 4).
+This highlights that the CF score and attention contribution scores capture different but complementary aspects of cultural bias in model representations.
+We also observe narrow connections for cultures with low coverage, such as Assam, Azerbaijan, and Ethiopia (Section 6.1.1).
+Rather than indicating cultural erasure through dominance by other cultures, this pattern likely reflects limited representational coverage of culturally distinctive knowledge of LLMs.
+In this sense, low-resource cultures may appear less susceptible to cultural flattening.
+This interpretation aligns with Li et al.
+(2025), who report weaker memorization of symbols associated with low-resource cultures in LLMs.
+Further evidence for this explanation comes from downstream behaviour: models exhibit a reduced preference for biased options involving low-resource cultures (Table 1), with the differences being statistically significant (Table 7).
+Taken together, these findings suggest that weaker cultural biases toward low-resource cultures do not stem from improved fairness, but rather from insufficient cultural representation.
+Consequently, improving model performance for low-resource cultures may require approaches that prioritize knowledge acquisition and representational coverage, rather than bias mitigation alone.
 9 Conclusion
-In this work, we investigate the complex and often biased ways LLMs process cultural knowledge.
-We introduce Culturescope, a method that leverages activation patching to probe the internal mechanisms of LLMs, allowing us to analyse the cultural knowledge encoded within their layers.
-In addition, we quantify the phenomenon of ‘cultural flattening’, where LLMs represent less-documented cultures through the concepts of more dominant or geographically close ones, thereby erasing cultural nuances.
-We conduct our research across three distinct models using culturally-grounded QA datasets, moving beyond isolated extrinsic evaluations to examine the interaction between cultural knowledge within the models’ parametric space.
-Our analysis reveals that models over-attend Western-centric tokens internally, indicating internalized Western-dominance bias.
-The results for low-resource cultures suggest that they may not be as susceptible to cultural biases as other cultures.
+In this work, we study how cultural biases arise from the internal representations of LLMs, by investigating LLMs’ internal representations.
+We introduce Culturescope, a method that leverages activation patching to analyse the cultural knowledge encoded within their layers.
+In addition, we quantify the phenomenon of ‘cultural flattening’, where LLMs represent less-documented cultures through the knowledge of more dominant or geographically close ones, thereby erasing cultural nuances.
 Our findings suggest that future work should develop a tailored approach that considers the impact of bias and resource levels to improve LLMs’ cultural understanding.
 Limitations
-While our study provides new insights into cultural knowledge in LLMs, there exists limitations.
 Although a growing number of benchmarks aim to evaluate cultural knowledge in LLMs, few are suitable for our evaluation setup.
 To meaningfully compare the probed cultural knowledge across cultures, the datasets must maintain a consistent QA format across different cultures.
-In addition, to ensure that the evaluation captures how LLMs represent culture on a global scale, it is essential to include a geographically diverse set of cultures.
+Thus, our methods and analyses inherit the limitations from the benchmarks, such as, lack of fine-grained set of cultures, and 1-to-1 mapping of a culture and a country.
 However, we emphasize that our proposed method is model- and task-agnostic, and can be applied to any dataset that meets these requirements.
 Due to computational constraints, we are reporting results with  8B models, unable to conduct experiments on larger-scale models.
-In addition, our analysis does not cover all countries globally.
-This limitation is not due to methodological oversight but rather the lack of publicly available datasets that include culturally grounded QA data for many regions.
-As such, our findings are necessarily constrained by the scope of existing resources.
-Finally, our MCQ with hard negatives (§3.2.3) involves a degree of random sampling.
+Our definition of low-resource cultures relies on the approximation of the most-spoken language from the previous work Joshi et al.
+(2020).
+Finally, our MCQ with hard negatives (Sec.
+3.2.3) involves a degree of random sampling.
 While this introduces some variability, we consider it a reasonable trade-off given the prohibitive cost of exhaustively evaluating all possible negative combinations.
 We mitigate this by ensuring consistency across runs and focusing on aggregate trends rather than individual instances.
-Acknowledgements
-[…]
 Appendix A Dataset Details
+In our experiments, we utilize BLEnD Myung et al.
+(2024) and CAMeL-2 Naous and Xu (2025).
+BLEnD comprises 500 short-answer question-answer pairs for each culture, where the answers vary depending on the cultural or regional context.
+To reduce computational cost, we exclude North Korea and West Java, resulting in a final selection of 14 cultural groups from BLEnD.
+With CAMeL-2, an LLM is asked to extract an entity from a context collected from Arabic X/Twitter data according to the specified entity type in the input.
+We take the English partition and reduce the dataset to 14 countries to keep a similar culture distribution to the BLEnD dataset.
+Dataset details, including domains and countries covered by the datasets, can be found in Appendix A.
+We provide their brief data statistics and characteristics in Table 6.
+A.1 Culture Groups
+As mentioned in Sec.
+3.2, our work conducts experiments that focus on 14 countries classified in two dimensions.
+Specifically, with the taxonomy from Joshi et al.
+(2020) which classifies languages into six levels (0: very low-resource to 5: very high-resource), we simplify this into three groups: High (Level 5), Mid (Levels 3-4), and Low (Levels 0-2).
+We assign each culture a language resource level based on its most widely spoken language, provided by Wikipedia.
+For the region dimension, we group countries into six regions based on continents.
+We split Asia into three subregions, which leaves us six regional groups: North America, Europe, Africa, West Asia, South Asia, and East Asia.
+We compare three groups based on the level of language resource Joshi et al.
+(2020) and six groups based on the continental region.
+Table 3 shows the culture entities that correspond to each group.
+BLEnD-Resource BLEnD-Region
+High 2548 Europe 789
+North America 520
+Mid 1425 East Asia 588
+West Asia 516
+Low 1077 South Asia 592
+Africa 771
+Sum 5050 Sum 3776
+Table 4: Number of questions per group in Cultural MCQ with hard negatives
+Option type Mean $\pm$ s.d. 95% CI
+BLEnD-Region
+Gold answer 269.71 $\pm$ 16.49 [260.2, 279.23]
+Hard negative 269.71 $\pm$ 16.3 [260.3, 279.12]
+Others 539.43 $\pm$ 38.34 [517.3, 561.56]
+BLEnD-Resource
+Gold answer 360.71 $\pm$ 11.4 [354.13, 367.30]
+High 721.43 $\pm$ 31.55 [692.25, 750.61]
+Mid 1683.33 $\pm$ 52.5 [1553.9, 1812.78]
+Low 1262.5 $\pm$ 85.57 [1126.49, 1398.51]
+Table 5: Descriptive statistics (mean, standard deviation (s.d.), and 95% Confidence Interval (CI)) of the number of selected culture groups for each option type in BLEnD-Resource and BLEnD-Region.
 Dataset Task Number of Questions Domain List of Countries
@@
 Commonsense QA
-5726
+5,726
 Education, Food, Holidays/Celebration/Leisure,
@@
 Naous and Xu (2025)
-Extractive QA 1862 Locations, Beverage, Food, Sports
+Extractive QA 1,862 Locations, Beverage, Food, Sports
 Africa: Morocco, Algeria
@@
 Table 6: Details for the datasets
-In our experiments, we utilize BLEnD Myung et al. (2024) and CAMeL-2 Naous and Xu (2025). We provide their brief data statistics and characteristics in Table 4.
-A.1 Country Groups
-As mentioned in Section 3.2, our work conducts experiments that focus on 14 countries classified in two dimensions.
-We compare three groups based on the level of language resource Joshi et al.
-(2020) and six groups based on the continental region.
-Table 3 shows the country entities that correspond to each group.
+BLEnD-Resource BLEnD-Region
+Hypothesis 1 Hypothesis 2 Hypothesis 3 Hypothesis 1 Hypothesis 2 Hypothesis 3
+Llama-3.1 low mid avg. 0.008 low low mid avg. 0.0049 low
+avg. 0.0034 0.6328 mid 0.2256 avg. 0.1066 avg. 0.0004 0.931 mid 0.0016 avg. 0.7306
+mid 0.0107 - low 0.076 mid 0.1008 mid 0.0001 - low 0.8067 mid 0.8815
+aya-expanse low mid avg. 0.009 low low mid avg. 0.0124 low
+avg. 0.0 0.936 mid 0.1477 avg. 0.0946 avg. 0.0007 0.8161 mid 0.0086 avg. 0.4606
+mid 0.0 - low 0.2307 mid 0.0548 mid 0.0007 - low 0.5762 mid 0.6784
+Qwen2.5 low mid avg. 0.0 low low mid avg. 0.0 low
+avg. 0.026 0.461 mid 0.22 avg. 0.3283 avg. 0.0 0.9722 mid 0.0122 avg. 0.1162
+mid 0.0796 - low 0.0215 mid 0.1636 mid 0.0 - low 0.0307 mid 0.1076
+Table 7: P-values from two-proportions Z-tests with three different hypotheses. Bolded texts mean that the test supports the hypothesis.
+A.2 Cultural MCQ with hard negatives
+The creation of BLEnD-Resource and BLEnD-Region involves random selection among possible options (Sec.
+3.2.3).
+Table 4 presents descriptive statistics including mean, standard deviation, and 95% confidence interval of the number of selected culture groups for each option type.
+This demonstrates that options are sampled to achieve a balanced representation of cultural groups.
+A gold answer option type refers to the culture group from which the option originates.
+An option type of hard negative refers to a culture group from which an option selected as a hard negative originates.
+High, mid, and low denote the number of selected countries from three resource groups.
 Appendix B Example of Prompts
 In Figure 5, we present the prompt templates we use for each method to obtain Table 2.
-Appendix C CF Score Results
+Appendix C Details for Culturescope
+BLEnD CAMeL-2
+Llama-3.1 English 0.325 $\pm$ 0.01 0.239 $\pm$ 0.07
+Spanish 0.277 $\pm$ 0.07 -
+Irrelevant 0.149 $\pm$ 0.09 0.155 $\pm$ 0.08
+aya-expanse English 0.349 $\pm$ 0.1 0.242 $\pm$ 0.07
+Irrelevant 0.127 $\pm$ 0.07 0.141 $\pm$ 0.06
+Qwen-2.5 English 0.344 $\pm$ 0.11 0.227 $\pm$ 0.07
+Arabic - 0.173 $\pm$ 0.07
+Irrelevant 0.141 $\pm$ 0.09 0.148 $\pm$ 0.07
+Table 8: Average and standard deviations of the similarity scores from the filtering stage of the Culturescope. We round the average off to two decimal places.
+CAMeL-2
+Relevance Evaluation
+Llama-3.1 aya-expanse Qwen2.5
+Baseline 0.7126 0.7019 0.6675
+Cultural Prompting 0.6799 0.7282 0.7148
+CANDLE 0.5640 0.6955 0.6473
+CultureScope 0.6519 0.7169 0.6659
+Table 9: We present the relevance evaluation of Culturescope on CAMeL-2 in English. We highlight the best-performing method in bold and the second-best with an underline.
+low vs high low vs mid
+Llama-3.1 1.11e-09 0.159
+aya-expense 8.55e-10 9.69e-13
+Qwen2.5 1.75e-07 5.54e-03
+Table 10: p-value of Student’s t-test for each model on low-resource target samples from BLEnD-Resource, based on attention scores.
+Comparisons were conducted between low and mid, low and high, all from $C_{resource}$, which are all hard negative options.
+During the filtering step in the Culturescope, we use three different semantic search models for each language.
+For English, we use a DeBERTa based model, trained for semantic search [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2.
+For Arabic, we use the bert-base model finetuned for semantic search in Arabic [^2] https://huggingface.co/akhooli/Arabic-SBERT-100K.
+Lastly, for Spanish, we use the bert-base model finetuned for semantic search in Spanish [^3] https://huggingface.co/hiiamsid/sentence_similarity_spanish_es.
+In Table 8, we show the threshold values selected for each dataset and language for the filtering stage.
+The values are selected by the average of cosine similarity score across all the generated concepts.
+Figure 6 describes the average number of generated knowledge per layer. This shows that Culturescope successfully generates relevant cultural knowledge from layers, especially from early layers.
+Appendix D Statistical Significance Tests
+D.1 Testing on MCQ Results
+To evaluate the statistical significance of our findings in Sec.
+6.2.1, we set three alternative hypotheses and test with a two-proportions Z-test  [^4] https://www.statsmodels.org/dev/generated/statsmodels.stats.proportion.proportions_ztest.html.
+Table 7 presents p-values from the list of Z-tests.
+avg., mid, and low follow the definition from Sec.
+6.2.1.
+The Z-test supports the alternative hypothesis when the p-value is lower than 0.05.
+Hypothesis 1 is that the accuracy from the avg.
+and the mid is higher than the accuracy from the low.
+The p-values in all models on both datasets support this hypothesis.
+Hypothesis 2 is that in each level, when an LLM predicts incorrectly, it prefers the biased option over the other options.
+This hypothesis is supported in avg.
+and mid groups, which highlights our finding on less susceptibility of low-resource cultures with cultural biases.
+Lastly, hypothesis 3 is that the model predicts biased options more in the low group than in the avg.
+and mid groups.
+This hypothesis is rejected in all tests, confirming the statistical significance of lower cultural bias in the low group.
+Llama-3.1 aya-expanse Qwen2.5
+Target Target vs EUR Target vs AME Target vs EUR Target vs AME Target vs EUR Target vs AME
+S-AS 0.0570 0.199 0.614 0.596 0.00109 0.0413
+E-AS 0.864 0.403 0.0851 0.787 0.799 0.609
+W-AS 0.827 0.476 0.787 0.829 0.329 0.475
+EUR 4.22e-03 0.216 2.26e-06 1.89e-04 0.160 0.104
+Table 11: p-value of Student’s t-test for each model on non-Western targeted samples from BLEnD-Region, based on attention scores.
+Comparisons were made between each sample’s target region and America, and between each sample’s target region and Europe, all within $C_{region}$, which represents hard negative options.
+D.2 Testing on the Results for Attention Contribution Score
+As discussed in Sec.
+6.1.2, our analyses suggest that high-resource bias and Western-dominance bias are strongly internalized within LLM representations, as visualized in the attention heatmaps in Figure 4 and Figure 8.
+To further validate these observations, we conduct Student’s t-tests [^5] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.ttest_ind.html on the attention scores of BLEnD-Resource and BLEnD-Region.
+To examine bias patterns, we conduct p-value analyses on attention scores using Student’s t-tests.
+For high-resource bias, low-resource targets ($c_{low}$) are compared against mid- and high-resource ($c_{mid}$, $c_{high}$) samples within $C_{resource}$, with most p-values less than 0.05, indicating statistically significant differences (Table 10).
+For Western-dominance bias, for non-Western regions (South Asia, East Asia, West Asia) targeted samples, non-Western regions ($c_{sameregion}$) are compared against Western regions (Europe, America) ($c_{region1}$, $c_{region2}$) within $C_{region}$.
+While fewer comparisons yield significance, several regions with large mean differences show p-values less than 0.05 (Table 11), suggesting Western-dominance bias is present but less uniformly across regions.
+In summary, the results confirm that high-resource bias is consistently strong in BLEnD-Resource, while Western-dominance bias is also evident in BLEnD-Region, though less uniformly across all regions.
+Appendix E CF Score Results
 In Figure 7, we show the CF score results with Llama-3.1, aya-expanse, Qwen2.5 on BLEnD and CAMeL-2.
-As mentioned in §6.2, we exclude the countries with CF scores lower than the average CF score across all countries.
-(a) BLEnD, aya-expanse
-(b) BLEnD, Qwen2.5
-(c) CAMeL-2, aya-expanse
-(d) CAMeL-2, Qwen2.5
-(e) CAMeL-2, Llama-3.1
+As mentioned in Sec.
+6.1.1, we exclude the countries with CF scores lower than the average CF score across all countries.
+1. Inspection prompts
+Generate associated words, Syria, Oman, Jordan, Qatar, West Asia, Turkey, Israel, Lebanon, …, Leonardo DiCaprio, Tom Cruise, Kate Winslet, Brad Pitt, Actor, …, Samsung, Cell Phone, TV, Apple, Nokia, South Korea, Electronics, …, x
+2.Open-ended QA Prompts
+2-1. Baseline
+BLEnD
+[…]
+BLEnD
+[…]
+BLEnD
+[…]
+Figure 5: Prompt templates used for open-ended QA evaluations.
+(a) BLEnD, English
+(b) CAMeL-2, English
+Figure 6: Average number of generated knowledge per layer.
+(a) BLEnD, English, aya-expanse
+(b) BLEnD, English, Qwen2.5
+(c) BLEnD, Spanish, Llama-3.1
+(d) CAMeL-2, English, Llama-3.1
+(e) CAMeL-2, English, aya-expanse
+(f) CAMeL-2, English, Qwen2.5
+(g) CAMeL-2, Arabic, Qwen2.5
 Figure 7: CF Score Results
-Appendix D Attention Contributions
+Appendix F Attention Contributions
 In cases where an option consists of multiple tokens, we follow the approach of Yuksekgonul et al.
-(2024), taking the maximum attention contribution score among the component tokens.
-When the attention scores are averaged by samples to examine its general patterns, simply averaging can be sensitive to extreme values or samples in which attention scores are globally biased toward higher or lower magnitudes.
+(2024) and take the maximum attention contribution score across the component tokens.
+When the attention scores are averaged by samples to examine their general patterns, simply averaging can be sensitive to extreme values or samples in which attention scores are globally biased toward higher or lower magnitudes.
 This variability can potentially reduce the generalizability of the results.